Autoconf

Creating Automatic Configuration Scripts

Edition 2.13, for Autoconf version 2.13 December 1998

by David MacKenzie and Ben Elliston

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 93, 94, 95, 96, 98 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

1 Introduction

A physicist, an engineer, and a computer scientist were discussing the nature of God. Surely a Physicist, said the physicist, because early in the Creation, God made Light; and you know, Maxwell’s equations, the dual nature of electro-magnetic waves, the relativist consequences… An Engineer!, said the engineer, because before making Light, God split the Chaos into Land and Water; it takes a hell of an engineer to handle that big amount of mud, and orderly separation of solids from liquids… The computer scientist shouted: And the Chaos, where do you think it was coming from, hmm?

—Anonymous Autoconf is a tool for producing shell scripts that automatically configure software source

code packages to adapt to many kinds of UNIX-like systems. The configuration scripts produced by Autoconf are independent of Autoconf when they are run, so their users do not need to have Autoconf.

The configuration scripts produced by Autoconf require no manual user intervention when run; they do not normally even need an argument specifying the system type. Instead, they test for the presence of each feature that the software package they are for might need individually. (Before each check, they print a one-line message stating what they are checking for, so the user doesn’t get too bored while waiting for the script to finish.) As a result, they deal well with systems that are hybrids or customized from the more common UNIX variants. There is no need to maintain files that list the features supported by each release of each variant of UNIX.

For each software package that Autoconf is used with, it creates a configuration script from a template file that lists the system features that the package needs or can use. After the shell code to recognize and respond to a system feature has been written, Autoconf allows it to be shared by many software packages that can use (or need) that feature. If it later turns out that the shell code needs adjustment for some reason, it needs to be changed in only one place; all of the configuration scripts can be regenerated automatically to take advantage of the updated code.

The Metaconfig package is similar in purpose to Autoconf, but the scripts it produces require manual user intervention, which is quite inconvenient when configuring large source trees. Unlike Metaconfig scripts, Autoconf scripts can support cross-compiling, if some care is taken in writing them.

There are several jobs related to making portable software packages that Autoconf currently does not do. Among these are automatically creating ‘Makefile’ files with all of the standard targets, and supplying replacements for standard library functions and header files on systems that lack them. Work is in progress to add those features in the future.

Autoconf imposes some restrictions on the names of macros used with #ifdef in C programs (see [Preprocessor Symbol Index], page 86).

Autoconf requires GNU m4 in order to generate the scripts. It uses features that some UNIX versions of m4 do not have. It also overflows internal limits of some versions of m4, including GNU m4 1.0. You must use version 1.1 or later of GNU m4. Using version 1.3 or later will be much faster than 1.1 or 1.2.

See Chapter 13 [Upgrading], page 71, for information about upgrading from version 1. See Chapter 14 [History], page 75, for the story of Autoconf’s development. See Chapter 12 [Questions], page 68, for answers to some common questions about Autoconf.

Mail suggestions and bug reports for Autoconf to bug-gnu-utils@prep.ai.mit.edu. Please include the Autoconf version number, which you can get by running ‘autoconf

--version’.

2 Making configure Scripts

The configuration scripts that Autoconf produces are by convention called configure. When run, configure creates several files, replacing configuration parameters in them with appropriate values. The files that configure creates are:

• one or more ‘Makefile’ files, one in each subdirectory of the package (see Section 3.3 [Makefile Substitutions], page 11);

• optionally, a C header file, the name of which is configurable, containing #define directives (see Section 3.4 [Configuration Headers], page 15);

• a shell script called ‘config.status’ that, when run, will recreate the files listed above (see ⟨undefined⟩ [Invoking config.status], page ⟨undefined⟩);

• a shell script called ‘config.cache’ that saves the results of running many of the tests (see Section 6.3.2 [Cache Files], page 49);

• a file called ‘config.log’ containing any messages produced by compilers, to help debugging if configure makes a mistake.

To create a configure script with Autoconf, you need to write an Autoconf input file ‘configure.in’ and run autoconf on it. If you write your own feature tests to supplement those that come with Autoconf, you might also write files called ‘aclocal.m4’ and ‘acsite.m4’. If you use a C header file to contain #define directives, you might also write ‘acconfig.h’, and you will distribute the Autoconf-generated file ‘config.h.in’ with the package.

Here is a diagram showing how the files that can be used in configuration are produced. Programs that are executed are suffixed by ‘*’. Optional files are enclosed in square brackets (‘[]’). autoconf and autoheader also read the installed Autoconf macro files (by reading

‘autoconf.m4’).

Files used in preparing a software package for distribution:

     your source files --> [autoscan*] --> [configure.scan] --> configure.in
     configure.in --.   .------> autoconf* -----> configure
                    +---+
     [aclocal.m4] --+   `---.
     [acsite.m4] ---'       |
                            +--> [autoheader*] -> [config.h.in]
     [acconfig.h] ----.     |
                      +-----'
     [config.h.top] --+
     [config.h.bot] --'
     Makefile.in -------------------------------> Makefile.in

Files used in configuring a software package:

                            .-------------> config.cache
     configure* ------------+-------------> config.log
                            |
     [config.h.in] -.       v            .-> [config.h] -.
                    +--> config.status* -+               +--> make*
     Makefile.in ---'                    `-> Makefile ---'
2.1 Writing ‘configure.in’

To produce a configure script for a software package, create a file called ‘configure.in’ that contains invocations of the Autoconf macros that test the system features your package needs or can use. Autoconf macros already exist to check for many features; see Chapter 4 [Existing Tests], page 19, for their descriptions. For most other features, you can use Autoconf template macros to produce custom checks; see Chapter 5 [Writing Tests], page 38, for information about them. For especially tricky or specialized features, ‘configure.in’ might need to contain some hand-crafted shell commands. The autoscan program can give you a good start in writing ‘configure.in’ (see Section 2.2 [Invoking autoscan], page 5, for more information).

The order in which ‘configure.in’ calls the Autoconf macros is not important, with a few exceptions. Every ‘configure.in’ must contain a call to AC_INIT before the checks, and a call to AC_OUTPUT at the end (see Section 3.2 [Output], page 9). Additionally, some macros rely on other macros having been called first, because they check previously set values of some variables to decide what to do. These macros are noted in the individual descriptions (see Chapter 4 [Existing Tests], page 19), and they also warn you when creating configure if they are called out of order.

To encourage consistency, here is a suggested order for calling the Autoconf macros. Generally speaking, the things near the end of this list could depend on things earlier in it. For example, library functions could be affected by typedefs and libraries.

     AC_INIT(file)

checks for programs checks for libraries checks for header files checks for typedefs checks for structures checks for compiler characteristics checks for library functions checks for system services

     AC_OUTPUT([file. . .])

It is best to put each macro call on its own line in ‘configure.in’. Most of the macros don’t add extra newlines; they rely on the newline after the macro call to terminate the commands. This approach makes the generated configure script a little easier to read by not inserting lots of blank lines. It is generally safe to set shell variables on the same line as a macro call, because the shell allows assignments without intervening newlines.

When calling macros that take arguments, there must not be any blank space between the macro name and the open parenthesis. Arguments can be more than one line long if they are enclosed within the m4 quote characters ‘[’ and ‘]’. If you have a long line such as a list of file names, you can generally use a backslash at the end of a line to continue it logically on the next line (this is implemented by the shell, not by anything special that Autoconf does).

Some macros handle two cases: what to do if the given condition is met, and what to do if the condition is not met. In some places you might want to do something if a condition is true but do nothing if it’s false, or vice versa. To omit the true case, pass an empty value for the action-if-found argument to the macro. To omit the false case, omit the action-if-not-found argument to the macro, including the comma before it.

You can include comments in ‘configure.in’ files by starting them with the m4 builtin macro dnl, which discards text up through the next newline. These comments do not appear in the generated configure scripts. For example, it is helpful to begin ‘configure.in’ files with a line like this:

     dnl Process this file with autoconf to produce a configure script.
2.2 Using autoscan to Create ‘configure.in’

The autoscan program can help you create a ‘configure.in’ file for a software package. autoscan examines source files in the directory tree rooted at a directory given as a command line argument, or the current directory if none is given. It searches the source files for common portability problems and creates a file ‘configure.scan’ which is a preliminary ‘configure.in’ for that package.

You should manually examine ‘configure.scan’ before renaming it to ‘configure.in’; it will probably need some adjustments. Occasionally autoscan outputs a macro in the wrong order relative to another macro, so that autoconf produces a warning; you need to move such macros manually. Also, if you want the package to use a configuration header file, you must add a call to AC_CONFIG_HEADER (see Section 3.4 [Configuration Headers], page 15). You might also have to change or add some #if directives to your program in order to make it work with Autoconf (see Section 2.3 [Invoking ifnames], page 6, for information about a program that can help with that job).

autoscan uses several data files, which are installed along with the distributed Autoconf macro files, to determine which macros to output when it finds particular symbols in a package’s source files. These files all have the same format. Each line consists of a symbol, whitespace, and the Autoconf macro to output if that symbol is encountered. Lines starting with ‘#’ are comments.

autoscan is only installed if you already have Perl installed. autoscan accepts the following options:

--help Print a summary of the command line options and exit.

--macrodir=dir

Look for the data files in directory dir instead of the default installation directory. You can also set the AC_MACRODIR environment variable to a directory; this option overrides the environment variable.

--verbose

Print the names of the files it examines and the potentially interesting symbols it finds in them. This output can be voluminous.

--version

Print the version number of Autoconf and exit.

2.3 Using ifnames to List Conditionals

ifnames can help when writing a ‘configure.in’ for a software package. It prints the identifiers that the package already uses in C preprocessor conditionals. If a package has already been set up to have some portability, this program can help you figure out what its configure needs to check for. It may help fill in some gaps in a ‘configure.in’ generated by autoscan (see Section 2.2 [Invoking autoscan], page 5).

ifnames scans all of the C source files named on the command line (or the standard input, if none are given) and writes to the standard output a sorted list of all the identifiers that appear in those files in #if, #elif, #ifdef, or #ifndef directives. It prints each identifier on a line, followed by a space-separated list of the files in which that identifier occurs. ifnames accepts the following options:

--help

-h Print a summary of the command line options and exit.

--macrodir=dir

-m dir Look for the Autoconf macro files in directory dir instead of the default installation directory. Only used to get the version number. You can also set the AC_MACRODIR environment variable to a directory; this option overrides the environment variable.

--version

Print the version number of Autoconf and exit.

2.4 Using autoconf to Create configure

To create configure from ‘configure.in’, run the autoconf program with no arguments. autoconf processes ‘configure.in’ with the m4 macro processor, using the Autoconf macros. If you give autoconf an argument, it reads that file instead of ‘configure.in’ and writes the configuration script to the standard output instead of to configure. If you give autoconf the argument ‘-’, it reads the standard input instead of ‘configure.in’ and writes the configuration script on the standard output.

The Autoconf macros are defined in several files. Some of the files are distributed with Autoconf; autoconf reads them first. Then it looks for the optional file ‘acsite.m4’ in the directory that contains the distributed Autoconf macro files, and for the optional file ‘aclocal.m4’ in the current directory. Those files can contain your site’s or the package’s own Autoconf macro definitions (see Chapter 7 [Writing Macros], page 52, for more information). If a macro is defined in more than one of the files that autoconf reads, the last definition it reads overrides the earlier ones.

autoconf accepts the following options:

--help

-h Print a summary of the command line options and exit.

--localdir=dir

-l dir Look for the package file ‘aclocal.m4’ in directory dir instead of in the current directory.

--macrodir=dir

-m dir Look for the installed macro files in directory dir. You can also set the AC_ MACRODIR environment variable to a directory; this option overrides the environment variable.

--version

Print the version number of Autoconf and exit.

2.5 Using autoreconf to Update configure Scripts

If you have a lot of Autoconf-generated configure scripts, the autoreconf program can save you some work. It runs autoconf (and autoheader, where appropriate) repeatedly to remake the Autoconf configure scripts and configuration header templates in the directory tree rooted at the current directory. By default, it only remakes those files that are older than their ‘configure.in’ or (if present) ‘aclocal.m4’. Since autoheader does not change the timestamp of its output file if the file wouldn’t be changing, this is not necessarily the minimum amount of work. If you install a new version of Autoconf, you can make autoreconf remake all of the files by giving it the ‘--force’ option.

If you give autoreconf the ‘--macrodir=dir’ or ‘--localdir=dir’ options, it passes them down to autoconf and autoheader (with relative paths adjusted properly).

autoreconf does not support having, in the same directory tree, both directories that are parts of a larger package (sharing ‘aclocal.m4’ and ‘acconfig.h’), and directories that are independent packages (each with their own ‘aclocal.m4’ and ‘acconfig.h’). It assumes that they are all part of the same package, if you use ‘--localdir’, or that each directory is a separate package, if you don’t use it. This restriction may be removed in the future.

See Section 3.3.3 [Automatic Remaking], page 14, for ‘Makefile’ rules to automatically remake configure scripts when their source files change. That method handles the timestamps of configuration header templates properly, but does not pass ‘--macrodir=dir’ or

‘--localdir=dir’.

autoreconf accepts the following options:

--help

-h Print a summary of the command line options and exit.

--force

-f Remake even ‘configure’ scripts and configuration headers that are newer than their input files (‘configure.in’ and, if present, ‘aclocal.m4’).

--localdir=dir

-l dir Have autoconf and autoheader look for the package files ‘aclocal.m4’ and (autoheader only) ‘acconfig.h’ (but not ‘file.top’ and ‘file.bot’) in directory dir instead of in the directory containing each ‘configure.in’.

--macrodir=dir

-m dir Look for the Autoconf macro files in directory dir instead of the default installation directory. You can also set the AC_MACRODIR environment variable to a directory; this option overrides the environment variable.

--verbose

Print the name of each directory where autoreconf runs autoconf (and autoheader, if appropriate).

--version

Print the version number of Autoconf and exit.

3 Initialization and Output Files

Autoconf-generated configure scripts need some information about how to initialize, such as how to find the package’s source files; and about the output files to produce. The following sections describe initialization and creating output files.

3.1 Finding configure Input

Every configure script must call AC_INIT before doing anything else. The only other required macro is AC_OUTPUT (see Section 3.2 [Output], page 9).

Macro AC INIT (unique-file-in-source-dir)

Process any command-line arguments and find the source code directory. unique-file-in-source-dir is some file that is in the package’s source directory; configure checks for this file’s existence to make sure that the directory that it is told contains the source code in fact does. Occasionally people accidentally specify the wrong directory with ‘--srcdir’; this is a safety check. See Chapter 10 [Invoking configure], page 65, for more information.

Packages that do manual configuration or use the install program might need to tell configure where to find some other shell scripts by calling AC_CONFIG_AUX_DIR, though the default places it looks are correct for most cases.

Macro AC CONFIG AUX DIR(dir)

Use the ‘install-sh’, ‘config.sub’, ‘config.guess’, and Cygnus configure

scripts that are in directory dir. These are auxiliary files used in configuration. dir can be either absolute or relative to ‘srcdir’. The default is ‘srcdir’ or ‘srcdir/..’ or ‘srcdir/../..’, whichever is the first that contains ‘install-sh’. The other files are not checked for, so that using AC_PROG_ INSTALL does not automatically require distributing the other auxiliary files. It checks for ‘install.sh’ also, but that name is obsolete because some make programs have a rule that creates ‘install’ from it if there is no ‘Makefile’.

3.2 Creating Output Files

Every Autoconf-generated configure script must finish by calling AC_OUTPUT. It is the macro that creates the ‘Makefile’s and optional other files resulting from configuration. The only other required macro is AC_INIT (see Section 3.1 [Input], page 9).

Macro AC OUTPUT ([file… [, extra-cmds [, init-cmds]]])

Create output files. Call this macro once, at the end of ‘configure.in’. The file… argument is a whitespace-separated list of output files; it may be empty. This macro creates each file ‘file’ by copying an input file (by default named ‘file.in’), substituting the output variable values. See Section 3.3 [Makefile Substitutions], page 11, for more information on using output variables. See Section 6.2 [Setting Output Variables], page 47, for more information on creating them. This macro creates the directory that the file is in if it doesn’t exist (but not the parents of that directory). Usually, ‘Makefile’s are created this way, but other files, such as ‘.gdbinit’, can be specified as well.

If AC_CONFIG_HEADER, AC_LINK_FILES, or AC_CONFIG_SUBDIRS has been called,

this macro also creates the files named as their arguments.

A typical call to AC_OUTPUT looks like this:

     AC_OUTPUT(Makefile src/Makefile man/Makefile X/Imakefile)

You can override an input file name by appending to file a colon-separated list of input files. Examples:

     AC_OUTPUT(Makefile:templates/top.mk lib/Makefile:templates/lib.mk)
     AC_OUTPUT(Makefile:templates/vars.mk:Makefile.in:templates/rules.mk)

Doing this allows you to keep your file names acceptable to MS-DOS, or to prepend and/or append boilerplate to the file.

If you pass extra-cmds, those commands will be inserted into ‘config.status’ to be run after all its other processing. If init-cmds are given, they are inserted just before extra-cmds, with shell variable, command, and backslash substitutions performed on them in configure. You can use init-cmds to pass variables from configure to the extra-cmds. If AC_OUTPUT_COMMANDS has been called, the commands given to it are run just before the commands passed to this macro.

Macro AC OUTPUT COMMANDS (extra-cmds [, init-cmds])

Specify additional shell commands to run at the end of ‘config.status’, and shell commands to initialize any variables from configure. This macro may be called multiple times. Here is an unrealistic example:

     fubar=27
     AC_OUTPUT_COMMANDS([echo this is extra $fubar, and so on.], fubar=$fubar)
     AC_OUTPUT_COMMANDS([echo this is another, extra, bit], [echo init bit])

If you run make on subdirectories, you should run it using the make variable MAKE. Most versions of make set MAKE to the name of the make program plus any options it was given. (But many do not include in it the values of any variables set on the command line, so those are not passed on automatically.) Some old versions of make do not set this variable. The following macro allows you to use it even with those versions.

Macro AC PROG MAKE SET

If make predefines the variable MAKE, define output variable SET_MAKE to be empty. Otherwise, define SET_MAKE to contain ‘MAKE=make’. Calls AC_SUBST

for SET_MAKE.

To use this macro, place a line like this in each ‘Makefile.in’ that runs MAKE on other directories:

@SET_MAKE@

3.3 Substitutions in Makefiles

Each subdirectory in a distribution that contains something to be compiled or installed should come with a file ‘Makefile.in’, from which configure will create a ‘Makefile’ in that directory. To create a ‘Makefile’, configure performs a simple variable substitution, replacing occurrences of ‘@variable@’ in ‘Makefile.in’ with the value that configure has determined for that variable. Variables that are substituted into output files in this way are called output variables. They are ordinary shell variables that are set in configure. To make configure substitute a particular variable into the output files, the macro AC_SUBST must be called with that variable name as an argument. Any occurrences of ‘@variable@’ for other variables are left unchanged. See Section 6.2 [Setting Output Variables], page 47, for more information on creating output variables with AC_SUBST.

A software package that uses a configure script should be distributed with a file ‘Makefile.in’, but no ‘Makefile’; that way, the user has to properly configure the package for the local system before compiling it.

See section “Makefile Conventions” in The GNU Coding Standards, for more information on what to put in ‘Makefile’s.

3.3.1 Preset Output Variables

Some output variables are preset by the Autoconf macros. Some of the Autoconf macros set additional output variables, which are mentioned in the descriptions for those macros. See [Output Variable Index], page 84, for a complete list of output variables. Here is what each of the preset ones contains. See section “Variables for Installation Directories” in The GNU Coding Standards, for more information about the variables with names that end in

‘dir’.

Variable bindir

The directory for installing executables that users run.

Variable configure input

A comment saying that the file was generated automatically by configure and giving the name of the input file. AC_OUTPUT adds a comment line containing this variable to the top of every ‘Makefile’ it creates. For other files, you should reference this variable in a comment at the top of each input file. For example, an input shell script should begin like this:

          #! /bin/sh
          # @configure_input@

The presence of that line also reminds people editing the file that it needs to be processed by configure in order to be used.

Variable datadir

The directory for installing read-only architecture-independent data.

Variable exec prefix

The installation prefix for architecture-dependent files.

Variable includedir

The directory for installing C header files.

Variable infodir

The directory for installing documentation in Info format.

Variable libdir

The directory for installing object code libraries.

Variable libexecdir

The directory for installing executables that other programs run.

Variable localstatedir

The directory for installing modifiable single-machine data.

Variable mandir

The top-level directory for installing documentation in man format.

Variable oldincludedir

The directory for installing C header files for non-gcc compilers.

Variable prefix

The installation prefix for architecture-independent files.

Variable sbindir

The directory for installing executables that system administrators run.

Variable sharedstatedir

The directory for installing modifiable architecture-independent data.

Variable srcdir

The directory that contains the source code for that ‘Makefile’.

Variable sysconfdir

The directory for installing read-only single-machine data.

Variable top srcdir

The top-level source code directory for the package. In the top-level directory, this is the same as srcdir.

Variable CFLAGS

Debugging and optimization options for the C compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_ PROG_CC (or empty if you don’t). configure uses this variable when compiling programs to test for C features.

Variable CPPFLAGS

Header file search directory (‘-Idir’) and any other miscellaneous options for the C preprocessor and compiler. If it is not set in the environment when configure runs, the default value is empty. configure uses this variable when compiling or preprocessing programs to test for C features.

Variable CXXFLAGS

Debugging and optimization options for the C++ compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_ PROG_CXX (or empty if you don’t). configure uses this variable when compiling programs to test for C++ features.

Variable FFLAGS

Debugging and optimization options for the Fortran 77 compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_PROG_F77 (or empty if you don’t). configure uses this variable when compiling programs to test for Fortran 77 features.

Variable DEFS

‘-D’ options to pass to the C compiler. If AC_CONFIG_HEADER is called, configure replaces ‘@DEFS@’ with ‘-DHAVE_CONFIG_H’ instead (see Section 3.4 [Configuration Headers], page 15). This variable is not defined while configure is performing its tests, only when creating the output files. See Section 6.2 [Setting Output Variables], page 47, for how to check the results of previous tests.

Variable LDFLAGS

Stripping (‘-s’) and any other miscellaneous options for the linker. If it is not set in the environment when configure runs, the default value is empty. configure uses this variable when linking programs to test for C features.

Variable LIBS

‘-l’ and ‘-L’ options to pass to the linker.

3.3.2 Build Directories

You can support compiling a software package for several architectures simultaneously from the same copy of the source code. The object files for each architecture are kept in their own directory.

To support doing this, make uses the VPATH variable to find the files that are in the source directory. GNU make and most other recent make programs can do this. Older make programs do not support VPATH; when using them, the source code must be in the same directory as the object files.

To support VPATH, each ‘Makefile.in’ should contain two lines that look like:

srcdir = @srcdir@
VPATH = @srcdir@

Do not set VPATH to the value of another variable, for example ‘VPATH = $(srcdir)’, because some versions of make do not do variable substitutions on the value of VPATH.

configure substitutes in the correct value for srcdir when it produces ‘Makefile’. Do not use the make variable $<, which expands to the pathname of the file in the source

directory (found with VPATH), except in implicit rules. (An implicit rule is one such as ‘.c.o’, which tells how to create a ‘.o’ file from a ‘.c’ file.) Some versions of make do not set $< in explicit rules; they expand it to an empty value.

Instead, ‘Makefile’ command lines should always refer to source files by prefixing them with ‘$(srcdir)/’. For example:

     time.info: time.texinfo
             $(MAKEINFO) $(srcdir)/time.texinfo

3.3.3 Automatic Remaking

You can put rules like the following in the top-level ‘Makefile.in’ for a package to automatically update the configuration information when you change the configuration files. This example includes all of the optional files, such as ‘aclocal.m4’ and those related to configuration header files. Omit from the ‘Makefile.in’ rules any of these files that your package does not use.

The ‘${srcdir}/’ prefix is included because of limitations in the VPATH mechanism.

The ‘stamp-’ files are necessary because the timestamps of ‘config.h.in’ and ‘config.h’ will not be changed if remaking them does not change their contents. This feature avoids unnecessary recompilation. You should include the file ‘stamp-h.in’ your package’s distribution, so make will consider ‘config.h.in’ up to date. On some old BSD systems, touch or any command that results in an empty file does not update the timestamps, so use a command like echo as a workaround.

     ${srcdir}/configure: configure.in aclocal.m4
             cd ${srcdir} && autoconf
     # autoheader might not change config.h.in, so touch a stamp file.
     ${srcdir}/config.h.in: stamp-h.in
     ${srcdir}/stamp-h.in: configure.in aclocal.m4 acconfig.h \
         config.h.top config.h.bot
             cd ${srcdir} && autoheader
             echo timestamp > ${srcdir}/stamp-h.in
     config.h: stamp-h
     stamp-h: config.h.in config.status
             ./config.status
     Makefile: Makefile.in config.status
             ./config.status
     config.status: configure
             ./config.status --recheck

In addition, you should pass ‘echo timestamp > stamp-h’ in the extra-cmds argument to AC_OUTPUT, so ‘config.status’ will ensure that ‘config.h’ is considered up to date. See Section 3.2 [Output], page 9, for more information about AC_OUTPUT.

See ⟨undefined⟩ [Invoking config.status], page ⟨undefined⟩, for more examples of handling configuration-related dependencies.

3.4 Configuration Header Files

When a package tests more than a few C preprocessor symbols, the command lines to pass ‘-D’ options to the compiler can get quite long. This causes two problems. One is that the make output is hard to visually scan for errors. More seriously, the command lines can exceed the length limits of some operating systems. As an alternative to passing ‘-D’ options to the compiler, configure scripts can create a C header file containing ‘#define’ directives. The AC_CONFIG_HEADER macro selects this kind of output. It should be called right after AC_INIT.

The package should ‘#include’ the configuration header file before any other header files, to prevent inconsistencies in declarations (for example, if it redefines const). Use

‘#include <config.h>’ instead of ‘#include "config.h"’, and pass the C compiler a ‘-I.’

option (or ‘-I..’; whichever directory contains ‘config.h’). That way, even if the source directory is configured itself (perhaps to make a distribution), other build directories can also be configured without finding the ‘config.h’ from the source directory.

Macro AC CONFIG HEADER (header-to-create …)

Make AC_OUTPUT create the file(s) in the whitespace-separated list header-to-create containing C preprocessor #define statements, and replace ‘@DEFS@’ in generated files with ‘-DHAVE_CONFIG_H’ instead of the value of DEFS. The usual name for header-to-create is ‘config.h’. If header-to-create already exists and its contents are identical to what AC_ OUTPUT would put in it, it is left alone. Doing this allows some changes in configuration without needlessly causing object files that depend on the header file to be recompiled. Usually the input file is named ‘header-to-create.in’; however, you can override the input file name by appending to header-to-create, a colon-separated list of input files. Examples:

          AC_CONFIG_HEADER(defines.h:defines.hin)
          AC_CONFIG_HEADER(defines.h:defs.pre:defines.h.in:defs.post)

Doing this allows you to keep your file names acceptable to MS-DOS, or to prepend and/or append boilerplate to the file.

3.4.1 Configuration Header Templates

Your distribution should contain a template file that looks as you want the final header file to look, including comments, with default values in the #define statements. For example, suppose your ‘configure.in’ makes these calls:

     AC_CONFIG_HEADER(conf.h)
     AC_CHECK_HEADERS(unistd.h)

Then you could have code like the following in ‘conf.h.in’. On systems that have ‘unistd.h’, configure will change the 0 to a 1. On other systems, it will leave the line unchanged.

     /* Define as 1 if you have unistd.h.  */
     #define HAVE_UNISTD_H 0

Alternately, if your code tests for configuration options using #ifdef instead of #if, a default value can be to #undef the variable instead of to define it to a value. On systems that

have ‘unistd.h’, configure will change the second line to read ‘#define HAVE_UNISTD_H

1’. On other systems, it will comment that line out (in case the system predefines that symbol).

     /* Define if you have unistd.h.  */
     #undef HAVE_UNISTD_H
3.4.2 Using autoheader to Create ‘config.h.in’

The autoheader program can create a template file of C ‘#define’ statements for

configure to use. If ‘configure.in’ invokes AC_CONFIG_HEADER(file), autoheader creates

‘file.in’; if multiple file arguments are given, the first one is used. Otherwise, autoheader

creates ‘config.h.in’.

If you give autoheader an argument, it uses that file instead of ‘configure.in’ and writes the header file to the standard output instead of to ‘config.h.in’. If you give autoheader an argument of ‘-’, it reads the standard input instead of ‘configure.in’ and writes the header file to the standard output.

autoheader scans ‘configure.in’ and figures out which C preprocessor symbols it might define. It copies comments and #define and #undef statements from a file called ‘acconfig.h’, which comes with and is installed with Autoconf. It also uses a file called ‘acconfig.h’ in the current directory, if present. If you AC_DEFINE any additional symbols, you must create that file with entries for them. For symbols defined by AC_CHECK_HEADERS,

AC_CHECK_FUNCS, AC_CHECK_SIZEOF, or AC_CHECK_LIB, autoheader generates comments

and #undef statements itself rather than copying them from a file, since the possible symbols are effectively limitless.

The file that autoheader creates contains mainly #define and #undef statements and their accompanying comments. If ‘./acconfig.h’ contains the string ‘@TOP@’, autoheader copies the lines before the line containing ‘@TOP@’ into the top of the file that it generates. Similarly, if ‘./acconfig.h’ contains the string ‘@BOTTOM@’, autoheader copies the lines after that line to the end of the file it generates. Either or both of those strings may be omitted.

An alternate way to produce the same effect is to create the files ‘file.top’ (typically ‘config.h.top’) and/or ‘file.bot’ in the current directory. If they exist, autoheader copies them to the beginning and end, respectively, of its output. Their use is discouraged because they have file names that contain two periods, and so can not be stored on MS-DOS; also, they are two more files to clutter up the directory. But if you use the ‘--localdir=dir’ option to use an ‘acconfig.h’ in another directory, they give you a way to put custom boilerplate in each individual ‘config.h.in’.

autoheader accepts the following options:

--help

-h Print a summary of the command line options and exit.

--localdir=dir

-l dir Look for the package files ‘aclocal.m4’ and ‘acconfig.h’ (but not ‘file.top’ and ‘file.bot’) in directory dir instead of in the current directory.

--macrodir=dir

-m dir Look for the installed macro files and ‘acconfig.h’ in directory dir. You can also set the AC_MACRODIR environment variable to a directory; this option overrides the environment variable.

--version

Print the version number of Autoconf and exit.

3.5 Configuring Other Packages in Subdirectories

In most situations, calling AC_OUTPUT is sufficient to produce ‘Makefile’s in subdirectories. However, configure scripts that control more than one independent package can use AC_CONFIG_SUBDIRS to run configure scripts for other packages in subdirectories.

Macro AC CONFIG SUBDIRS (dir …)

Make AC_OUTPUT run configure in each subdirectory dir in the given whitespaceseparated list. If a given dir is not found, no error is reported, so a configure script can configure whichever parts of a large source tree are present. If a given dir contains ‘configure.in’ but no configure, the Cygnus configure script found by AC_CONFIG_AUXDIR is used. The subdirectory configure scripts are given the same command line options that were given to this configure script, with minor changes if needed (e.g., to adjust a relative path for the cache file or source directory). This macro also sets the output variable subdirs to the list of directories ‘dir …’. ‘Makefile’ rules can use this variable to determine which subdirectories to recurse into. This macro may be called multiple times.

3.6 Default Prefix

By default, configure sets the prefix for files it installs to ‘/usr/local’. The user of configure can select a different prefix using the ‘--prefix’ and ‘--exec-prefix’ options. There are two ways to change the default: when creating configure, and when running it.

Some software packages might want to install in a directory besides ‘/usr/local’ by default. To accomplish that, use the AC_PREFIX_DEFAULT macro.

Macro AC PREFIX DEFAULT (prefix)

Set the default installation prefix to prefix instead of ‘/usr/local’.

It may be convenient for users to have configure guess the installation prefix from the location of a related program that they have already installed. If you wish to do that, you

can call AC_PREFIX_PROGRAM.

Macro AC PREFIX PROGRAM (program)

If the user did not specify an installation prefix (using the ‘--prefix’ option), guess a value for it by looking for program in PATH, the way the shell does. If program is found, set the prefix to the parent of the directory containing program; otherwise leave the prefix specified in ‘Makefile.in’ unchanged. For example, if program is gcc and the PATH contains ‘/usr/local/gnu/bin/gcc’, set the prefix to ‘/usr/local/gnu’.

3.7 Version Numbers in configure

The following macros manage version numbers for configure scripts. Using them is optional.

Macro AC PREREQ (version)

Ensure that a recent enough version of Autoconf is being used. If the version of Autoconf being used to create configure is earlier than version, print an error message on the standard error output and do not create configure. For example:

     AC_PREREQ(1.8)

This macro is useful if your ‘configure.in’ relies on non-obvious behavior that changed between Autoconf releases. If it merely needs recently added macros, then AC_PREREQ is less useful, because the autoconf program already tells the user which macros are not found. The same thing happens if ‘configure.in’ is processed by a version of Autoconf older than when AC_PREREQ was added.

Macro AC REVISION (revision-info)

Copy revision stamp revision-info into the configure script, with any dollar signs or double-quotes removed. This macro lets you put a revision stamp from ‘configure.in’ into configure without RCS or CVS changing it when you check in configure. That way, you can determine easily which revision of ‘configure.in’ a particular configure corresponds to. It is a good idea to call this macro before AC_INIT so that the revision number is near the top of both ‘configure.in’ and configure. To support doing that, the AC_REVISION output begins with ‘#! /bin/sh’, like the normal start of a configure script does. For example, this line in ‘configure.in’:

     AC_REVISION($Revision: 1.30 $)dnl

produces this in configure:

     #! /bin/sh
     # From configure.in Revision: 1.30

4 Existing Tests

These macros test for particular system features that packages might need or want to use. If you need to test for a kind of feature that none of these macros check for, you can probably do it by calling primitive test macros with appropriate arguments (see Chapter 5 [Writing Tests], page 38).

These tests print messages telling the user which feature they’re checking for, and what they find. They cache their results for future configure runs (see Section 6.3 [Caching Results], page 48).

Some of these macros set output variables. See Section 3.3 [Makefile Substitutions], page 11, for how to get their values. The phrase “define name” is used below as a shorthand to mean “define C preprocessor symbol name to the value 1”. See Section 6.1 [Defining Symbols], page 46, for how to get those symbol definitions into your program.

4.1 Alternative Programs

These macros check for the presence or behavior of particular programs. They are used to choose between several alternative programs and to decide what to do once one has been chosen. If there is no macro specifically defined to check for a program you need, and you don’t need to check for any special properties of it, then you can use one of the general program check macros.

4.1.1 Particular Program Checks

These macros check for particular programs—whether they exist, and in some cases whether they support certain features.

Macro AC DECL YYTEXT

Define YYTEXT_POINTER if yytext is a ‘char *’ instead of a ‘char []’. Also set output variable LEX_OUTPUT_ROOT to the base of the file name that the lexer generates; usually ‘lex.yy’, but sometimes something else. These results vary according to whether lex or flex is being used.

Macro AC PROG AWK

Check for mawk, gawk, nawk, and awk, in that order, and set output variable AWK to the first one that it finds. It tries mawk first because that is reported to be the fastest implementation.

Macro AC PROG CC

Determine a C compiler to use. If CC is not already set in the environment, check for gcc, and use cc if that’s not found. Set output variable CC to the name of the compiler found. If using the GNU C compiler, set shell variable GCC to ‘yes’, empty otherwise. If output variable CFLAGS was not already set, set it to ‘-g -O2’ for the GNU C compiler (‘-O2’ on systems where GCC does not accept ‘-g’), or ‘-g’ for other compilers.

If the C compiler being used does not produce executables that can run on the system where configure is being run, set the shell variable cross_compiling to ‘yes’, otherwise ‘no’. In other words, this tests whether the build system type is different from the host system type (the target system type is irrelevant to this test). See Chapter 8 [Manual Configuration], page 56, for more on support for cross compiling.

Macro AC PROG CC C O

If the C compiler does not accept the ‘-c’ and ‘-o’ options simultaneously,

define NO_MINUS_C_MINUS_O.

Macro AC PROG CPP

Set output variable CPP to a command that runs the C preprocessor. If ‘$CC -E’ doesn’t work, it uses ‘/lib/cpp’. It is only portable to run CPP on files with a ‘.c’ extension. If the current language is C (see Section 5.8 [Language Choice], page 44), many of the specific test macros use the value of CPP indirectly by calling AC_TRY_CPP,

AC_CHECK_HEADER, AC_EGREP_HEADER, or AC_EGREP_CPP.

Macro AC PROG CXX

Determine a C++ compiler to use. Check if the environment variable CXX or CCC (in that order) is set; if so, set output variable CXX to its value. Otherwise search for a C++ compiler under likely names (c++, g++, gcc, CC, cxx, and cc++). If none of those checks succeed, as a last resort set CXX to gcc. If using the GNU C++ compiler, set shell variable GXX to ‘yes’, empty otherwise. If output variable CXXFLAGS was not already set, set it to ‘-g -O2’ for the GNU C++ compiler (‘-O2’ on systems where G++ does not accept ‘-g’), or ‘-g’ for other compilers. If the C++ compiler being used does not produce executables that can run on the system where configure is being run, set the shell variable cross_compiling to ‘yes’, otherwise ‘no’. In other words, this tests whether the build system type is different from the host system type (the target system type is irrelevant to this test). See Chapter 8 [Manual Configuration], page 56, for more on support for cross compiling.

Macro AC PROG CXXCPP

Set output variable CXXCPP to a command that runs the C++ preprocessor. If ‘$CXX -E’ doesn’t work, it uses ‘/lib/cpp’. It is only portable to run CXXCPP on files with a ‘.c’, ‘.C’, or ‘.cc’ extension. If the current language is C++ (see Section 5.8 [Language Choice], page 44), many of the specific test macros use the value of CXXCPP indirectly by calling

AC_TRY_CPP, AC_CHECK_HEADER, AC_EGREP_HEADER, or AC_EGREP_CPP.

Macro AC PROG F77

Determine a Fortran 77 compiler to use. If F77 is not already set in the environment, check for g77, f77 and f2c, in that order. Set the output variable F77 to the name of the compiler found.

If using g77 (the GNU Fortran 77 compiler), then AC_PROG_F77 will set the shell variable G77 to ‘yes’, and empty otherwise. If the output variable FFLAGS was not already set in the environment, then set it to ‘-g -02’ for g77 (or ‘-O2’ where g77 does not accept ‘-g’). Otherwise, set FFLAGS to ‘-g’ for all other Fortran 77 compilers.

Macro AC PROG F77 C O

Test if the Fortran 77 compiler accepts the options ‘-c’ and ‘-o’ simultaneously,

and define F77_NO_MINUS_C_MINUS_O if it does not.

Macro AC PROG GCC TRADITIONAL

Add ‘-traditional’ to output variable CC if using the GNU C compiler and ioctl does not work properly without ‘-traditional’. That usually happens when the fixed header files have not been installed on an old system. Since recent versions of the GNU C compiler fix the header files automatically when installed, this is becoming a less prevalent problem.

Macro AC PROG INSTALL

Set output variable INSTALL to the path of a BSD compatible install program, if one is found in the current PATH. Otherwise, set INSTALL to ‘dir/install-sh -c’, checking the directories specified to AC_CONFIG_AUX_DIR (or its default directories) to determine dir (see Section 3.2 [Output], page 9). Also set the vari-

ables INSTALL_PROGRAM and INSTALL_SCRIPT to ‘${INSTALL}’ and INSTALL_
DATA to ‘${INSTALL} -m 644’.

This macro screens out various instances of install known to not work. It prefers to find a C program rather than a shell script, for speed. Instead of ‘install-sh’, it can also use ‘install.sh’, but that name is obsolete because some make programs have a rule that creates ‘install’ from it if there is no

‘Makefile’.

A copy of ‘install-sh’ which you may use comes with Autoconf. If you use

AC_PROG_INSTALL, you must include either ‘install-sh’ or ‘install.sh’ in

your distribution, or configure will produce an error message saying it can’t find them—even if the system you’re on has a good install program. This check is a safety measure to prevent you from accidentally leaving that file out, which would prevent your package from installing on systems that don’t have a BSD-compatible install program. If you need to use your own installation program because it has features not found in standard install programs, there is no reason to use AC_PROG_ INSTALL; just put the pathname of your program into your ‘Makefile.in’ files.

Macro AC PROG LEX

If flex is found, set output variable LEX to ‘flex’ and LEXLIB to ‘-lfl’, if that library is in a standard place. Otherwise set LEX to ‘lex’ and LEXLIB to ‘-ll’.

Macro AC PROG LN S

If ‘ln -s’ works on the current filesystem (the operating system and filesystem support symbolic links), set output variable LN_S to ‘ln -s’, otherwise set it to ‘ln’.

If the link is put in a directory other than the current directory, its meaning depends on whether ‘ln’ or ‘ln -s’ is used. To safely create links using ‘$(LN_S)’, either find out which form is used and adjust the arguments, or always invoke ln in the directory where the link is to be created. In other words, it does not work to do

     $(LN_S) foo /x/bar

Instead, do

     (cd /x && $(LN_S) foo bar)

Macro AC PROG RANLIB

Set output variable RANLIB to ‘ranlib’ if ranlib is found, otherwise to ‘:’ (do nothing).

Macro AC PROG YACC

If bison is found, set output variable YACC to ‘bison -y’. Otherwise, if byacc is found, set YACC to ‘byacc’. Otherwise set YACC to ‘yacc’.

4.1.2 Generic Program and File Checks

These macros are used to find programs not covered by the particular test macros. If you need to check the behavior of a program as well as find out whether it is present, you have to write your own test for it (see Chapter 5 [Writing Tests], page 38). By default, these macros use the environment variable PATH. If you need to check for a program that might not be in the user’s PATH, you can pass a modified path to use instead, like this:

AC_PATH_PROG(INETD, inetd, /usr/libexec/inetd,
  $PATH:/usr/libexec:/usr/sbin:/usr/etc:etc)

Macro AC CHECK FILE (file [, action-if-found [, action-if-not-found]])

Check whether file file exists on the native system. If it is found, execute action-if-found, otherwise do action-if-not-found, if given.

Macro AC CHECK FILES (files[, action-if-found [, action-if-not-found]])

Executes AC_CHECK_FILE once for each file listed in files. Additionally, defines ‘HAVEfile’ for each file found, set to 1.

Macro AC CHECK PROG (variable, prog-to-check-for, value-if-found [,

value-if-not-found [, path, [ reject ]]]) Check whether program prog-to-check-for exists in PATH. If it is found, set variable to value-if-found, otherwise to value-if-not-found, if given. Always pass over reject (an absolute file name) even if it is the first found in the search path; in that case, set variable using the absolute file name of the prog-to-check-for found that is not reject. If variable was already set, do nothing. Calls AC_SUBST for variable.

Macro AC CHECK PROGS (variable, progs-to-check-for [,

value-if-not-found [, path]]) Check for each program in the whitespace-separated list progs-to-check-for exists in PATH. If it is found, set variable to the name of that program. Otherwise, continue checking the next program in the list. If none of the programs in the list are found, set variable to value-if-not-found; if value-if-not-found is not specified, the value of variable is not changed. Calls AC_SUBST for variable.

Macro AC CHECK TOOL (variable, prog-to-check-for [,

value-if-not-found [, path]]) Like AC_CHECK_PROG, but first looks for prog-to-check-for with a prefix of the host type as determined by AC_CANONICAL_HOST, followed by a dash (see Section 8.2 [Canonicalizing], page 56). For example, if the user runs ‘configure

--host=i386-gnu’, then this call:
     AC_CHECK_TOOL(RANLIB, ranlib, :)

sets RANLIB to ‘i386-gnu-ranlib’ if that program exists in PATH, or to ‘ranlib’ if that program exists in PATH, or to ‘:’ if neither program exists.

Macro AC PATH PROG (variable, prog-to-check-for [, value-if-not-found

[, path]]) Like AC_CHECK_PROG, but set variable to the entire path of prog-to-check-for if found.

Macro AC PATH PROGS (variable, progs-to-check-for [,

value-if-not-found [, path]]) Like AC_CHECK_PROGS, but if any of progs-to-check-for are found, set variable to the entire path of the program found.

4.2 Library Files

The following macros check for the presence of certain C, C++ or Fortran 77 library archive files.

Macro AC CHECK LIB (library, function [, action-if-found [,

action-if-not-found [, other-libraries]]]) Depending on the current language(see Section 5.8 [Language Choice], page 44), try to ensure that the C, C++ or Fortran 77 function function is available by checking whether a test program can be linked with the library library to get the function. library is the base name of the library; e.g., to check for ‘-lmp’, use ‘mp’ as the library argument.

action-if-found is a list of shell commands to run if the link with the library succeeds; action-if-not-found is a list of shell commands to run if the link fails. If action-if-found is not specified, the default action will add ‘-llibrary’ to LIBS and define ‘HAVE_LIBlibrary’ (in all capitals).

If linking with library results in unresolved symbols, which would be resolved by linking with additional libraries, give those libraries as the other-libraries argument, separated by spaces: ‘-lXt -lX11’. Otherwise this macro will fail to detect that library is present, because linking the test program will always fail with unresolved symbols.

Macro AC HAVE LIBRARY (library, [, action-if-found [,

action-if-not-found [, other-libraries]]]) This macro is equivalent to calling AC_CHECK_LIB with a function argument of main. In addition, library can be written as any of ‘foo’, ‘-lfoo’, or ‘libfoo.a’. In all of those cases, the compiler is passed ‘-lfoo’. However, library can not be a shell variable; it must be a literal name. This macro is considered obsolete.

Macro AC SEARCH LIBS (function, search-libs [, action-if-found [,

action-if-not-found [, other-libraries]]]) Search for a library defining function, if it’s not already available. This equates to calling AC_TRY_LINK_FUNC first with no libraries, then for each library listed in search-libs. If the function is found, run action-if-found, otherwise run action-if-not-found. If linking with library results in unresolved symbols, which would be resolved by linking with additional libraries, give those libraries as the other-libraries argument, separated by spaces: ‘-lXt -lX11’. Otherwise this macro will fail to detect that function is present, because linking the test program will always fail with unresolved symbols.

Macro AC SEARCH LIBS (function, search-libs[, action-if-found [,

action-if-not-found]]) This macro is equivalent to calling AC_TRY_LINK_FUNC once for each library listed in search-libs. Add ‘-llibrary’ to LIBS for the first library found to contain function, and execute action-if-found. Otherwise execute action-if-not-found.

4.3 Library Functions

The following macros check for particular C library functions. If there is no macro specifically defined to check for a function you need, and you don’t need to check for any special properties of it, then you can use one of the general function check macros.

4.3.1 Particular Function Checks

These macros check for particular C functions—whether they exist, and in some cases how they respond when given certain arguments.

Macro AC FUNC ALLOCA

Check how to get alloca. Tries to get a builtin version by checking for ‘alloca.h’ or the predefined C preprocessor macros __GNUC__and _AIX. If this macro finds ‘alloca.h’, it defines HAVE_ALLOCA_H. If those attempts fail, it looks for the function in the standard C library. If any of those methods succeed, it defines HAVE_ALLOCA. Otherwise, it sets the output variable ALLOCA to ‘alloca.o’ and defines C_ALLOCA (so programs can periodically call ‘alloca(0)’ to garbage collect). This variable is separate from LIBOBJS so multiple programs can share the value of ALLOCA without needing to create an actual library, in case only some of them use the code in LIBOBJS.

This macro does not try to get alloca from the System V R3 ‘libPW’ or the System V R4 ‘libucb’ because those libraries contain some incompatible functions that cause trouble. Some versions do not even contain alloca or contain a buggy version. If you still want to use their alloca, use ar to extract ‘alloca.o’ from them instead of compiling ‘alloca.c’.

Source files that use alloca should start with a piece of code like the following, to declare it properly. In some versions of AIX, the declaration of alloca must precede everything else except for comments and preprocessor directives. The #pragma directive is indented so that pre-ANSI C compilers will ignore it, rather than choke on it.

     /* AIX requires this to be the first thing in the file.  */
     #ifndef __GNUC__
     # if HAVE_ALLOCA_H
     #  include <alloca.h>
     # else
     #  ifdef _AIX
      #pragma alloca
     #  else
     #   ifndef alloca /* predefined by HP cc +Olibcalls */
     char *alloca ();
     #   endif
     #  endif
     # endif
     #endif

Macro AC FUNC CLOSEDIR VOID

If the closedir function does not return a meaningful value, define CLOSEDIR_ VOID. Otherwise, callers ought to check its return value for an error indicator.

Macro AC FUNC FNMATCH

If the fnmatch function is available and works (unlike the one on SunOS 5.4),

define HAVE_FNMATCH.

Macro AC FUNC GETLOADAVG

Check how to get the system load averages. If the system has the getloadavg function, this macro defines HAVE_GETLOADAVG, and adds to LIBS any libraries needed to get that function.

Otherwise, it adds ‘getloadavg.o’ to the output variable LIBOBJS, and possibly defines several other C preprocessor macros and output variables:

1. It defines SVR4, DGUX, UMAX, or UMAX4_3 if on those systems.

2. If it finds ‘nlist.h’, it defines NLIST_STRUCT.

 3. If ‘struct nlist’ has an ‘n_un’ member, it defines NLIST_NAME_UNION.

4. If compiling ‘getloadavg.c’ defines LDAV_PRIVILEGED, programs need to be installed specially on this system for getloadavg to work, and this

    macro defines GETLOADAVG_PRIVILEGED.

5. This macro sets the output variable NEED_SETGID. The value is ‘true’ if special installation is required, ‘false’ if not. If NEED_SETGID is ‘true’, this macro sets KMEM_GROUP to the name of the group that should own the installed program.

Macro AC FUNC GETMNTENT

Check for getmntent in the ‘sun’, ‘seq’, and ‘gen’ libraries, for Irix 4, PTX, and Unixware, respectively. Then, if getmntent is available, define HAVE_

GETMNTENT.

Macro AC FUNC GETPGRP

If getpgrp takes no argument (the POSIX.1 version), define GETPGRP_VOID. Otherwise, it is the BSD version, which takes a process ID as an argument. This macro does not check whether getpgrp exists at all; if you need to work in that situation, first call AC_CHECK_FUNC for getpgrp.

Macro AC FUNC MEMCMP

If the memcmp function is not available, or does not work on 8-bit data (like the one on SunOS 4.1.3), add ‘memcmp.o’ to output variable LIBOBJS.

Macro AC FUNC MMAP

If the mmap function exists and works correctly, define HAVE_MMAP. Only checks private fixed mapping of already-mapped memory.

Macro AC FUNC SELECT ARGTYPES

Determines the correct type to be passed to each of the select function’s ar-

guments, and defines those types in SELECT_TYPE_ARG1, SELECT_TYPE_ARG234,
and SELECT_TYPE_ARG5 respectively. SELECT_TYPE_ARG1 defaults to ‘int’,
SELECT_TYPE_ARG234 defaults to ‘int *’, and SELECT_TYPE_ARG5 defaults to
‘struct timeval *’.

Macro AC FUNC SETPGRP

If setpgrp takes no argument (the POSIX.1 version), define SETPGRP_VOID. Otherwise, it is the BSD version, which takes two process ID as arguments. This macro does not check whether setpgrp exists at all; if you need to work in that situation, first call AC_CHECK_FUNC for setpgrp.

Macro AC FUNC SETVBUF REVERSED

If setvbuf takes the buffering type as its second argument and the buffer pointer as the third, instead of the other way around, define SETVBUF_REVERSED. This is the case on System V before release 3.

Macro AC FUNC STRCOLL

If the strcoll function exists and works correctly, define HAVE_STRCOLL. This does a bit more than ‘AC_CHECK_FUNCS(strcoll)’, because some systems have incorrect definitions of strcoll, which should not be used.

Macro AC FUNC STRFTIME

Check for strftime in the ‘intl’ library, for SCO UNIX. Then, if strftime is available, define HAVE_STRFTIME.

Macro AC FUNC UTIME NULL

If ‘utime(file, NULL)’ sets file’s timestamp to the present, define HAVE_UTIME_

NULL.

Macro AC FUNC VFORK

If ‘vfork.h’ is found, define HAVE_VFORK_H. If a working vfork is not found, define vfork to be fork. This macro checks for several known errors in implementations of vfork and considers the system to not have a working vfork if it detects any of them. It is not considered to be an implementation error if a child’s invocation of signal modifies the parent’s signal handler, since child processes rarely change their signal handlers.

Macro AC FUNC VPRINTF

If vprintf is found, define HAVE_VPRINTF. Otherwise, if _doprnt is found, define HAVE_DOPRNT. (If vprintf is available, you may assume that vfprintf and vsprintf are also available.)

Macro AC FUNC WAIT3

If wait3 is found and fills in the contents of its third argument (a ‘struct rusage *’), which HP-UX does not do, define HAVE_WAIT3.

4.3.2 Generic Function Checks

These macros are used to find functions not covered by the particular test macros. If the functions might be in libraries other than the default C library, first call AC_CHECK_LIB for those libraries. If you need to check the behavior of a function as well as find out whether it is present, you have to write your own test for it (see Chapter 5 [Writing Tests], page 38).

Macro AC CHECK FUNC (function, [action-if-found [,

action-if-not-found]]) If C function function is available, run shell commands action-if-found, otherwise action-if-not-found. If you just want to define a symbol if the function is available, consider using AC_CHECK_FUNCS instead. This macro checks for functions with C linkage even when AC_LANG_CPLUSPLUS has been called, since C++ is more standardized than C is. (see Section 5.8 [Language Choice], page 44, for more information about selecting the language for checks.)

Macro AC CHECK FUNCS (function… [, action-if-found [,

action-if-not-found]]) For each given function in the whitespace-separated argument list that is available, define HAVE_function (in all capitals). If action-if-found is given, it is additional shell code to execute when one of the functions is found. You can give it a value of ‘break’ to break out of the loop on the first match. If actionif-not-found is given, it is executed when one of the functions is not found.

Macro AC REPLACE FUNCS (function…)

Like calling AC_CHECK_FUNCS using an action-if-not-found that adds ‘function.o’ to the value of the output variable LIBOBJS. You can declare a function for which your replacement version is used by enclosing the prototype in ‘#ifndef HAVE_function’. If the system has the function, it probably declares it in a header file you should be including, so you shouldn’t redeclare it, lest your declaration conflict.

4.4 Header Files

The following macros check for the presence of certain C header files. If there is no macro specifically defined to check for a header file you need, and you don’t need to check for any special properties of it, then you can use one of the general header file check macros.

4.4.1 Particular Header Checks

These macros check for particular system header files—whether they exist, and in some cases whether they declare certain symbols.

Macro AC DECL SYS SIGLIST

Define SYS_SIGLIST_DECLARED if the variable sys_siglist is declared in a

system header file, either ‘signal.h’ or ‘unistd.h’.

Macro AC DIR HEADER

Like calling AC_HEADER_DIRENT and AC_FUNC_CLOSEDIR_VOID, but defines a

different set of C preprocessor macros to indicate which header file is found. This macro and the names it defines are considered obsolete. The names it defines are:

‘dirent.h’
          DIRENT
‘sys/ndir.h’
          SYSNDIR
‘sys/dir.h’
          SYSDIR
‘ndir.h’ NDIR

In addition, if the closedir function does not return a meaningful value, define

VOID_CLOSEDIR.

Macro AC HEADER DIRENT

Check for the following header files, and for the first one that is found and defines ‘DIR’, define the listed C preprocessor macro:

‘dirent.h’
          HAVE_DIRENT_H
‘sys/ndir.h’
          HAVE_SYS_NDIR_H
‘sys/dir.h’
          HAVE_SYS_DIR_H

‘ndir.h’ HAVE_NDIR_H

The directory library declarations in the source code should look something like the following:

     #if HAVE_DIRENT_H
     # include <dirent.h>
     # define NAMLEN(dirent) strlen((dirent)->d_name)
     #else
     # define dirent direct
     # define NAMLEN(dirent) (dirent)->d_namlen
     # if HAVE_SYS_NDIR_H
     #  include <sys/ndir.h>
     # endif
     # if HAVE_SYS_DIR_H
     #  include <sys/dir.h>
     # endif
     # if HAVE_NDIR_H
     #  include <ndir.h>
     # endif
     #endif

Using the above declarations, the program would declare variables to be type struct dirent, not struct direct, and would access the length of a directory entry name by passing a pointer to a struct dirent to the NAMLEN macro. This macro also checks for the SCO Xenix ‘dir’ and ‘x’ libraries.

Macro AC HEADER MAJOR

If ‘sys/types.h’ does not define major, minor, and makedev, but ‘sys/mkdev.h’

does, define MAJOR_IN_MKDEV; otherwise, if ‘sys/sysmacros.h’ does, define

MAJOR_IN_SYSMACROS.

Macro AC HEADER STDC

Define STDC_HEADERS if the system has ANSI C header files. Specifically, this macro checks for ‘stdlib.h’, ‘stdarg.h’, ‘string.h’, and ‘float.h’; if the system has those, it probably has the rest of the ANSI C header files. This macro also checks whether ‘string.h’ declares memchr (and thus presumably the other mem functions), whether ‘stdlib.h’ declare free (and thus presumably malloc and other related functions), and whether the ‘ctype.h’ macros work on characters with the high bit set, as ANSI C requires. Use STDC_HEADERS instead of __STDC__to determine whether the system has ANSI-compliant header files (and probably C library functions) because many systems that have GCC do not have ANSI C header files. On systems without ANSI C headers, there is so much variation that it is probably easier to declare the functions you use than to figure out exactly what the system header files declare. Some systems contain a mix of functions ANSI and BSD; some are mostly ANSI but lack ‘memmove’; some define the BSD functions as macros in ‘string.h’ or ‘strings.h’; some have only the BSD functions but ‘string.h’; some declare the memory functions in ‘memory.h’, some in ‘string.h’; etc. It is probably sufficient to check for one string function and one memory function; if the library has the ANSI versions of those then it probably has most of the others. If you put the following in ‘configure.in’:

AC_HEADER_STDC
AC_CHECK_FUNCS(strchr memcpy)

then, in your code, you can put declarations like this:

#if STDC_HEADERS
# include <string.h>
#else
# ifndef HAVE_STRCHR
#  define strchr index
#  define strrchr rindex
# endif
char *strchr (), *strrchr ();
# ifndef HAVE_MEMCPY
#  define memcpy(d, s, n) bcopy ((s), (d), (n))
#  define memmove(d, s, n) bcopy ((s), (d), (n))
# endif
#endif

If you use a function like memchr, memset, strtok, or strspn, which have no BSD equivalent, then macros won’t suffice; you must provide an implementation of each function. An easy way to incorporate your implementations only when needed (since the ones in system C libraries may be hand optimized) is to, taking memchr for example, put it in ‘memchr.c’ and use

‘AC_REPLACE_FUNCS(memchr)’.

Macro AC HEADER SYS WAIT

If ‘sys/wait.h’ exists and is compatible with POSIX.1, define HAVE_SYS_WAIT_ H. Incompatibility can occur if ‘sys/wait.h’ does not exist, or if it uses the old BSD union wait instead of int to store a status value. If ‘sys/wait.h’ is not POSIX.1 compatible, then instead of including it, define the POSIX.1 macros with their usual interpretations. Here is an example:

#include <sys/types.h>
#if HAVE_SYS_WAIT_H
# include <sys/wait.h>
#endif
#ifndef WEXITSTATUS
# define WEXITSTATUS(stat_val) ((unsigned)(stat_val) >> 8)
#endif
#ifndef WIFEXITED
# define WIFEXITED(stat_val) (((stat_val) & 255) == 0)
#endif

Macro AC MEMORY H

Define NEED_MEMORY_H if memcpy, memcmp, etc. are not declared in ‘string.h’

and ‘memory.h’ exists. This macro is obsolete; instead, use AC_CHECK_

HEADERS(memory.h). See the example for AC_HEADER_STDC.

Macro AC UNISTD H

Define HAVE_UNISTD_H if the system has ‘unistd.h’. This macro is obsolete;

instead, use ‘AC_CHECK_HEADERS(unistd.h)’.

The way to check if the system supports POSIX.1 is:

     #if HAVE_UNISTD_H
     # include <sys/types.h>
     # include <unistd.h>
     #endif
     #ifdef _POSIX_VERSION
     /* Code for POSIX.1 systems.  */
     #endif

_POSIX_VERSION is defined when ‘unistd.h’ is included on POSIX.1 systems. If there is no ‘unistd.h’, it is definitely not a POSIX.1 system. However, some non-POSIX.1 systems do have ‘unistd.h’.

Macro AC USG

Define USG if the system does not have ‘strings.h’, rindex, bzero, etc. This implies that it has ‘string.h’, strrchr, memset, etc. The symbol USG is obsolete. Instead of this macro, see the example for AC_

HEADER_STDC.

4.4.2 Generic Header Checks

These macros are used to find system header files not covered by the particular test macros. If you need to check the contents of a header as well as find out whether it is present, you have to write your own test for it (see Chapter 5 [Writing Tests], page 38).

Macro AC CHECK HEADER (header-file, [action-if-found [,

action-if-not-found]]) If the system header file header-file exists, execute shell commands action-if-found, otherwise execute action-if-not-found. If you just want to define a symbol if the header file is available, consider using AC_CHECK_HEADERS instead.

Macro AC CHECK HEADERS (header-file… [, action-if-found [,

action-if-not-found]]) For each given system header file header-file in the whitespace-separated argument list that exists, define HAVE_header-file (in all capitals). If action-if-found is given, it is additional shell code to execute when one of the header files is found. You can give it a value of ‘break’ to break out of the loop on the first match. If action-if-not-found is given, it is executed when one of the header files is not found.

4.5 Structures

The following macros check for certain structures or structure members. To check structures not listed here, use AC_EGREP_CPP (see Section 5.1 [Examining Declarations], page 38) or AC_TRY_COMPILE (see Section 5.2 [Examining Syntax], page 39).

Macro AC HEADER STAT

If the macros S_ISDIR, S_ISREG et al. defined in ‘sys/stat.h’ do not work properly (returning false positives), define STAT_MACROS_BROKEN. This is the case on Tektronix UTekV, Amdahl UTS and Motorola System V/88.

Macro AC HEADER TIME

If a program may include both ‘time.h’ and ‘sys/time.h’, define TIME_ WITH_SYS_TIME. On some older systems, ‘sys/time.h’ includes ‘time.h’, but ‘time.h’ is not protected against multiple inclusion, so programs should not explicitly include both files. This macro is useful in programs that use, for

example, struct timeval or struct timezone as well as struct tm. It is best

used in conjunction with HAVE_SYS_TIME_H, which can be checked for using

AC_CHECK_HEADERS(sys/time.h).
     #if TIME_WITH_SYS_TIME
     # include <sys/time.h>
     # include <time.h>
     #else
     # if HAVE_SYS_TIME_H
     #  include <sys/time.h>
     # else
     #  include <time.h>
     # endif
     #endif

Macro AC STRUCT ST BLKSIZE

If struct stat contains an st_blksize member, define HAVE_ST_BLKSIZE.

Macro AC STRUCT ST BLOCKS

If struct stat contains an st_blocks member, define HAVE_ST_BLOCKS. Oth-

erwise, add ‘fileblocks.o’ to the output variable LIBOBJS.

Macro AC STRUCT ST RDEV

If struct stat contains an st_rdev member, define HAVE_ST_RDEV.

Macro AC STRUCT TM

If ‘time.h’ does not define struct tm, define TM_IN_SYS_TIME, which means that including ‘sys/time.h’ had better define struct tm.

Macro AC STRUCT TIMEZONE

Figure out how to get the current timezone. If struct tm has a tm_zone member, define HAVE_TM_ZONE. Otherwise, if the external array tzname is found,

define HAVE_TZNAME.

4.6 Typedefs

The following macros check for C typedefs. If there is no macro specifically defined to check for a typedef you need, and you don’t need to check for any special properties of it, then you can use a general typedef check macro.

4.6.1 Particular Typedef Checks

These macros check for particular C typedefs in ‘sys/types.h’ and ‘stdlib.h’ (if it exists).

Macro AC TYPE GETGROUPS

Define GETGROUPS_T to be whichever of gid_t or int is the base type of the array argument to getgroups.

Macro AC TYPE MODE T

If mode_t is not defined, define mode_t to be int.

Macro AC TYPE OFF T

If off_t is not defined, define off_t to be long.

Macro AC TYPE PID T

If pid_t is not defined, define pid_t to be int.

Macro AC TYPE SIGNAL

If ‘signal.h’ declares signal as returning a pointer to a function returning void, define RETSIGTYPE to be void; otherwise, define it to be int. Define signal handlers as returning type RETSIGTYPE:

     RETSIGTYPE
     hup_handler ()
     {

…

     }

Macro AC TYPE SIZE T

If size_t is not defined, define size_t to be unsigned.

Macro AC TYPE UID T

If uid_t is not defined, define uid_t to be int and gid_t to be int.

4.6.2 Generic Typedef Checks

This macro is used to check for typedefs not covered by the particular test macros.

Macro AC CHECK TYPE (type, default)

If the type type is not defined in ‘sys/types.h’, or ‘stdlib.h’ or ‘stddef.h’ if they exist, define it to be the C (or C++) builtin type default; e.g., ‘short’

or ‘unsigned’.

4.7 C Compiler Characteristics

The following macros check for C compiler or machine architecture features. To check for characteristics not listed here, use AC_TRY_COMPILE (see Section 5.2 [Examining Syntax], page 39) or AC_TRY_RUN (see Section 5.4 [Run Time], page 41)

Macro AC C BIGENDIAN

If words are stored with the most significant byte first (like Motorola and SPARC, but not Intel and VAX, CPUs), define WORDS_BIGENDIAN.

Macro AC C CONST

If the C compiler does not fully support the keyword const, define const to be empty. Some C compilers that do not define __STDC__do support const; some compilers that define __STDC__do not completely support const. Programs can simply use const as if every C compiler supported it; for those that don’t, the ‘Makefile’ or configuration header file will define it as empty.

Macro AC C INLINE

If the C compiler supports the keyword inline, do nothing. Otherwise define inline to __inline__or __inline if it accepts one of those, otherwise define inline to be empty.

Macro AC C CHAR UNSIGNED

If the C type char is unsigned, define __CHAR_UNSIGNED__, unless the C compiler predefines it.

Macro AC C LONG DOUBLE

If the C compiler supports the long double type, define HAVE_LONG_DOUBLE. Some C compilers that do not define __STDC__do support the long double type; some compilers that define __STDC__do not support long double.

Macro AC C STRINGIZE

If the C preprocessor supports the stringizing operator, define HAVE_STRINGIZE. The stringizing operator is ‘#’ and is found in macros such as this:

     #define x(y) #y

Macro AC CHECK SIZEOF (type [, cross-size])

Define SIZEOF_uctype to be the size in bytes of the C (or C++) builtin type type, e.g. ‘int’ or ‘char *’. If ‘type’ is unknown to the compiler, it gets a size of 0. uctype is type, with lowercase converted to uppercase, spaces changed to underscores, and asterisks changed to ‘P’. If cross-compiling, the value cross-size is used if given, otherwise configure exits with an error message. For example, the call

     AC_CHECK_SIZEOF(int *)

defines SIZEOF_INT_P to be 8 on DEC Alpha AXP systems.

Macro AC INT 16 BITS

If the C type int is 16 bits wide, define INT_16_BITS. This macro is obsolete; it is more general to use ‘AC_CHECK_SIZEOF(int)’ instead.

Macro AC LONG 64 BITS

If the C type long int is 64 bits wide, define LONG_64_BITS. This macro is obsolete; it is more general to use ‘AC_CHECK_SIZEOF(long)’ instead.

4.8 Fortran 77 Compiler Characteristics

The following macros check for Fortran 77 compiler characteristics. To check for characteristics not listed here, use AC_TRY_COMPILE (see Section 5.2 [Examining Syntax], page 39) or AC_TRY_RUN (see Section 5.4 [Run Time], page 41), making sure to first set the current lanuage to Fortran 77 AC_LANG_FORTRAN77 (see Section 5.8 [Language Choice], page 44).

Macro AC F77 LIBRARY LDFLAGS

Determine the linker flags (e.g. ‘-L’ and ‘-l’) for the Fortran 77 intrinsic and run-time libraries that are required to successfully link a Fortran 77 program or shared library. The output variable FLIBS is set to these flags. This macro is intended to be used in those situations when it is necessary to mix, e.g. C++ and Fortran 77 source code into a single program or shared library (see section “Mixing Fortran 77 With C and C++” in GNU Automake). For example, if object files from a C++ and Fortran 77 compiler must be linked together, then the C++ compiler/linker must be used for linking (since special C++-ish things need to happen at link time like calling global constructors, instantiating templates, enabling exception support, etc.). However, the Fortran 77 intrinsic and run-time libraries must be linked in as well, but the C++ compiler/linker doesn’t know by default how to add these Fortran 77 libraries. Hence, the macro AC_F77_LIBRARY_LDFLAGS was created to determine these Fortran 77 libraries.

4.9 System Services

The following macros check for operating system services or capabilities.

Macro AC CYGWIN

Checks for the Cygwin environment. If present, sets shell variable CYGWIN to ‘yes’. If not present, sets CYGWIN to the empty string.

Macro AC EXEEXT

Defines substitute variable EXEEXT based on the output of the compiler, after .c, .o, and .obj files have been excluded. Typically set to empty string if Unix, ‘.exe’ or ‘.EXE’ if Win32.

Macro AC OBJEXT

Defines substitute variable OBJEXT based on the output of the compiler, after .c files have been excluded. Typically set to ‘.o’ if Unix, ‘.obj’ if Win32.

Macro AC MINGW32

Checks for the MingW32 compiler environment. If present, sets shell variable MINGW32 to ‘yes’. If not present, sets MINGW32 to the empty string.

Macro AC PATH X

Try to locate the X Window System include files and libraries. If the user gave the command line options ‘--x-includes=dir’ and ‘--x-libraries=dir’, use those directories. If either or both were not given, get the missing values by running xmkmf on a trivial ‘Imakefile’ and examining the ‘Makefile’ that it produces. If that fails (such as if xmkmf is not present), look for them in several directories where they often reside. If either method is successful, set the shell variables x_includes and x_libraries to their locations, unless they are in directories the compiler searches by default.

If both methods fail, or the user gave the command line option ‘--without-x’, set the shell variable no_x to ‘yes’; otherwise set it to the empty string.

Macro AC PATH XTRA

An enhanced version of AC_PATH_X. It adds the C compiler flags that X needs to output variable X_CFLAGS, and the X linker flags to X_LIBS. If X is not

available, adds ‘-DX_DISPLAY_MISSING’ to X_CFLAGS.

This macro also checks for special libraries that some systems need in order to compile X programs. It adds any that the system needs to output variable X_EXTRA_LIBS. And it checks for special X11R6 libraries that need to be linked with before ‘-lX11’, and adds any found to the output variable X_PRE_LIBS.

Macro AC SYS INTERPRETER

Check whether the system supports starting scripts with a line of the form ‘#! /bin/csh’ to select the interpreter to use for the script. After running this macro, shell code in configure.in can check the shell variable interpval; it will be set to ‘yes’ if the system supports ‘#!’, ‘no’ if not.

Macro AC SYS LONG FILE NAMES

If the system supports file names longer than 14 characters, define HAVE_LONG_

FILE_NAMES.

Macro AC SYS RESTARTABLE SYSCALLS

If the system automatically restarts a system call that is interrupted by a signal,

define HAVE_RESTARTABLE_SYSCALLS.

4.10 UNIX Variants

The following macros check for certain operating systems that need special treatment for some programs, due to exceptional oddities in their header files or libraries. These macros are warts; they will be replaced by a more systematic approach, based on the functions they make available or the environments they provide.

Macro AC AIX

If on AIX, define _ALL_SOURCE. Allows the use of some BSD functions. Should be called before any macros that run the C compiler.

Macro AC DYNIX SEQ

If on Dynix/PTX (Sequent UNIX), add ‘-lseq’ to output variable LIBS. This macro is obsolete; instead, use AC_FUNC_GETMNTENT.

Macro AC IRIX SUN

If on IRIX (Silicon Graphics UNIX), add ‘-lsun’ to output variable LIBS. This macro is obsolete. If you were using it to get getmntent, use AC_FUNC_ GETMNTENT instead. If you used it for the NIS versions of the password and

group functions, use ‘AC_CHECK_LIB(sun, getpwnam)’.

Macro AC ISC POSIX

If on a POSIXized ISC UNIX, define _POSIX_SOURCE and add ‘-posix’ (for the GNU C compiler) or ‘-Xp’ (for other C compilers) to output variable CC. This allows the use of POSIX facilities. Must be called after AC_PROG_CC and before any other macros that run the C compiler.

Macro AC MINIX

If on Minix, define _MINIX and _POSIX_SOURCE and define _POSIX_1_SOURCE

to be 2. This allows the use of POSIX facilities. Should be called before any macros that run the C compiler.

Macro AC SCO INTL

If on SCO UNIX, add ‘-lintl’ to output variable LIBS. This macro is obsolete;

instead, use AC_FUNC_STRFTIME.

Macro AC XENIX DIR

If on Xenix, add ‘-lx’ to output variable LIBS. Also, if ‘dirent.h’ is being used, add ‘-ldir’ to LIBS. This macro is obsolete; use AC_HEADER_DIRENT instead.

5 Writing Tests

If the existing feature tests don’t do something you need, you have to write new ones. These macros are the building blocks. They provide ways for other macros to check whether various kinds of features are available and report the results.

This chapter contains some suggestions and some of the reasons why the existing tests are written the way they are. You can also learn a lot about how to write Autoconf tests by looking at the existing ones. If something goes wrong in one or more of the Autoconf tests, this information can help you understand the assumptions behind them, which might help you figure out how to best solve the problem.

These macros check the output of the C compiler system. They do not cache the results of their tests for future use (see Section 6.3 [Caching Results], page 48), because they don’t know enough about the information they are checking for to generate a cache variable name. They also do not print any messages, for the same reason. The checks for particular kinds of C features call these macros and do cache their results and print messages about what they’re checking for.

When you write a feature test that could be applicable to more than one software package, the best thing to do is encapsulate it in a new macro. See Chapter 7 [Writing Macros], page 52, for how to do that.

5.1 Examining Declarations

The macro AC_TRY_CPP is used to check whether particular header files exist. You can check for one at a time, or more than one if you need several header files to all exist for some purpose.

Macro AC TRY CPP (includes, [action-if-true [, action-if-false]])

includes is C or C++#include statements and declarations, on which shell variable, backquote, and backslash substitutions are performed. (Actually, it can be any C program, but other statements are probably not useful.) If the preprocessor produces no error messages while processing it, run shell commands action-if-true. Otherwise run shell commands action-if-false. This macro uses CPPFLAGS, but not CFLAGS, because ‘-g’, ‘-O’, etc. are not valid options to many C preprocessors.

Here is how to find out whether a header file contains a particular declaration, such as a typedef, a structure, a structure member, or a function. Use AC_EGREP_HEADER instead of running grep directly on the header file; on some systems the symbol might be defined in another header file that the file you are checking ‘#include’s.

Macro AC EGREP HEADER (pattern, header-file, action-if-found [,

action-if-not-found]) If the output of running the preprocessor on the system header file header-file matches the egrep regular expression pattern, execute shell commands actionif-found, otherwise execute action-if-not-found.

To check for C preprocessor symbols, either defined by header files or predefined by the C preprocessor, use AC_EGREP_CPP. Here is an example of the latter:

AC_EGREP_CPP(yes,
[#ifdef _AIX
  yes
#endif
], is_aix=yes, is_aix=no)

Macro AC EGREP CPP (pattern, program, [action-if-found [,

action-if-not-found]]) program is the text of a C or C++ program, on which shell variable, backquote, and backslash substitutions are performed. If the output of running the preprocessor on program matches the egrep regular expression pattern, execute shell commands action-if-found, otherwise execute action-if-not-found. This macro calls AC_PROG_CPP or AC_PROG_CXXCPP (depending on which language is current, see Section 5.8 [Language Choice], page 44), if it hasn’t been called already.

5.2 Examining Syntax

To check for a syntax feature of the C, C++ or Fortran 77 compiler, such as whether it recognizes a certain keyword, use AC_TRY_COMPILE to try to compile a small program that uses that feature. You can also use it to check for structures and structure members that are not present on all systems.

Macro AC TRY COMPILE (includes, function-body, [action-if-found [,

action-if-not-found]]) Create a C, C++ or Fortran 77 test program (depending on which language is current, see Section 5.8 [Language Choice], page 44), to see whether a function whose body consists of function-body can be compiled. For C and C++, includes is any #include statements needed by the code in function-body (includes will be ignored if the currently selected language is Fortran 77). This macro also uses CFLAGS or CXXFLAGS if either C or C++ is the currently selected language, as well as CPPFLAGS, when compiling. If Fortran 77 is the currently selected language then FFLAGS will be used when compiling. If the file compiles successfully, run shell commands action-if-found, otherwise run action-if-not-found. This macro does not try to link; use AC_TRY_LINK if you need to do that (see Section 5.3 [Examining Libraries], page 39).

5.3 Examining Libraries

To check for a library, a function, or a global variable, Autoconf configure scripts try to compile and link a small program that uses it. This is unlike Metaconfig, which by default uses nm or ar on the C library to try to figure out which functions are available. Trying to link with the function is usually a more reliable approach because it avoids dealing with the variations in the options and output formats of nm and ar and in the location of the standard libraries. It also allows configuring for cross-compilation or checking a function’s runtime behavior if needed. On the other hand, it can be slower than scanning the libraries once.

A few systems have linkers that do not return a failure exit status when there are unresolved functions in the link. This bug makes the configuration scripts produced by Autoconf unusable on those systems. However, some of them can be given options that make the exit status correct. This is a problem that Autoconf does not currently handle automatically. If users encounter this problem, they might be able to solve it by setting LDFLAGS in the environment to pass whatever options the linker needs (for example, ‘-Wl,-dn’ on MIPS RISC/OS).

AC_TRY_LINK is used to compile test programs to test for functions and global variables. It is also used by AC_CHECK_LIB to check for libraries (see Section 4.2 [Libraries], page 23), by adding the library being checked for to LIBS temporarily and trying to link a small program.

Macro AC TRY LINK (includes, function-body, [action-if-found [,

action-if-not-found]]) Depending on the current language (see Section 5.8 [Language Choice], page 44), create a test program to see whether a function whose body consists of functionbody can be compiled and linked.

For C and C++, includes is any #include statements needed by the code in function-body (includes will be ignored if the currently selected language is Fortran 77). This macro also uses CFLAGS or CXXFLAGS if either C or C++ is the currently selected language, as well as CPPFLAGS, when compiling. If Fortran 77 is the currently selected language then FFLAGS will be used when compiling. However, both LDFLAGS and LIBS will be used during linking in all cases.

If the file compiles and links successfully, run shell commands action-if-found, otherwise run action-if-not-found.

Macro AC TRY LINK FUNC (function, [action-if-found [,

action-if-not-found]]) Depending on the current language (see Section 5.8 [Language Choice], page 44), create a test program to see whether a program whose body consists of a prototype of and a call to function can be compiled and linked.

If the file compiles and links successfully, run shell commands action-if-found, otherwise run action-if-not-found.

Macro AC TRY LINK FUNC (function, [action-if-found [,

action-if-not-found]]) Attempt to compile and link a small program that links with function. If the file compiles and links successfully, run shell commands action-if-found, otherwise run action-if-not-found.

Macro AC COMPILE CHECK (echo-text, includes, function-body,

action-if-found [, action-if-not-found]) This is an obsolete version of AC_TRY_LINK, with the addition that it prints ‘checking for echo-text’ to the standard output first, if echo-text is non-empty. Use AC_MSG_CHECKING and AC_MSG_RESULT instead to print messages (see Section 6.4 [Printing Messages], page 50).

5.4 Checking Run Time Behavior

Sometimes you need to find out how a system performs at run time, such as whether a given function has a certain capability or bug. If you can, make such checks when your program runs instead of when it is configured. You can check for things like the machine’s endianness when your program initializes itself.

If you really need to test for a run-time behavior while configuring, you can write a test program to determine the result, and compile and run it using AC_TRY_RUN. Avoid running test programs if possible, because using them prevents people from configuring your package for cross-compiling.

5.4.1 Running Test Programs

Use the following macro if you need to test run-time behavior of the system while configuring.

Macro AC TRY RUN (program, [action-if-true [, action-if-false [,

action-if-cross-compiling]]]) program is the text of a C program, on which shell variable and backquote substitutions are performed. If it compiles and links successfully and returns an exit status of 0 when executed, run shell commands action-if-true. Otherwise run shell commands action-if-false; the exit status of the program is available in the shell variable ‘$?’. This macro uses CFLAGS or CXXFLAGS, CPPFLAGS, LDFLAGS, and LIBS when compiling. If the C compiler being used does not produce executables that run on the system where configure is being run, then the test program is not run. If the optional shell commands action-if-cross-compiling are given, they are run instead. Otherwise, configure prints an error message and exits.

Try to provide a pessimistic default value to use when cross-compiling makes runtime tests impossible. You do this by passing the optional last argument to AC_TRY_RUN. autoconf prints a warning message when creating configure each time it encounters a call to AC_TRY_RUN with no action-if-cross-compiling argument given. You may ignore the warning, though users will not be able to configure your package for cross-compiling. A few of the macros distributed with Autoconf produce this warning message.

To configure for cross-compiling you can also choose a value for those parameters based on the canonical system name (see Chapter 8 [Manual Configuration], page 56). Alternatively, set up a test results cache file with the correct values for the target system (see Section 6.3 [Caching Results], page 48).

To provide a default for calls of AC_TRY_RUN that are embedded in other macros, including a few of the ones that come with Autoconf, you can call AC_PROG_CC before running them. Then, if the shell variable cross_compiling is set to ‘yes’, use an alternate method to get the results instead of calling the macros.

Macro AC C CROSS

This macro is obsolete; it does nothing.

5.4.2 Guidelines for Test Programs

Test programs should not write anything to the standard output. They should return 0 if the test succeeds, nonzero otherwise, so that success can be distinguished easily from a core dump or other failure; segmentation violations and other failures produce a nonzero exit status. Test programs should exit, not return, from main, because on some systems (old Suns, at least) the argument to return in main is ignored.

Test programs can use #if or #ifdef to check the values of preprocessor macros defined by tests that have already run. For example, if you call AC_HEADER_STDC, then later on in ‘configure.in’ you can have a test program that includes an ANSI C header file conditionally:

     #if STDC_HEADERS
     # include <stdlib.h>
     #endif

If a test program needs to use or create a data file, give it a name that starts with ‘conftest’, such as ‘conftestdata’. The configure script cleans up by running ‘rm -rf conftest*’ after running test programs and if the script is interrupted.

5.4.3 Test Functions

Function declarations in test programs should have a prototype conditionalized for C++. In practice, though, test programs rarely need functions that take arguments.

     #ifdef __cplusplus
     foo(int i)
     #else
     foo(i) int i;
     #endif

Functions that test programs declare should also be conditionalized for C++, which requires ‘extern "C"’ prototypes. Make sure to not include any header files containing clashing prototypes.

     #ifdef __cplusplus
     extern "C" void *malloc(size_t);
     #else
     char *malloc();
     #endif

If a test program calls a function with invalid parameters (just to see whether it exists), organize the program to ensure that it never invokes that function. You can do this by calling it in another function that is never invoked. You can’t do it by putting it after a call to exit, because GCC version 2 knows that exit never returns and optimizes out any code that follows it in the same block.

If you include any header files, make sure to call the functions relevant to them with the correct number of arguments, even if they are just 0, to avoid compilation errors due to prototypes. GCC version 2 has internal prototypes for several functions that it automatically inlines; for example, memcpy. To avoid errors when checking for them, either pass them the correct number of arguments or redeclare them with a different return type (such as char).

5.5 Portable Shell Programming

When writing your own checks, there are some shell script programming techniques you should avoid in order to make your code portable. The Bourne shell and upward-compatible shells like Bash and the Korn shell have evolved over the years, but to prevent trouble, do not take advantage of features that were added after UNIX version 7, circa 1977. You should not use shell functions, aliases, negated character classes, or other features that are not found in all Bourne-compatible shells; restrict yourself to the lowest common denominator. Even unset is not supported by all shells! Also, include a space after the exclamation point in interpreter specifications, like this:

     #! /usr/bin/perl

If you omit the space before the path, then 4.2BSD based systems (such as Sequent DYNIX) will ignore the line, because they interpret ‘#! /’ as a 4-byte magic number.

The set of external programs you should run in a configure script is fairly small. See section “Utilities in Makefiles” in GNU Coding Standards, for the list. This restriction allows users to start out with a fairly small set of programs and build the rest, avoiding too many interdependencies between packages.

Some of these external utilities have a portable subset of features, as well; for example, don’t rely on ln having a ‘-f’ option or cat having any options. sed scripts should not contain comments or use branch labels longer than 8 characters. Don’t use ‘grep -s’ to suppress output, because ‘grep -s’ on System V does not suppress output, only error messages. Instead, redirect the standard output and standard error (in case the file doesn’t exist) of grep to ‘/dev/null’. Check the exit status of grep to determine whether it found a match.

5.6 Testing Values and Files

configure scripts need to test properties of many files and strings. Here are some portability problems to watch out for when doing those tests.

The test program is the way to perform many file and string tests. It is often invoked by the alternate name ‘[’, but using that name in Autoconf code is asking for trouble since it is an m4 quote character.

If you need to make multiple checks using test, combine them with the shell operators ‘&&’ and ‘||’ instead of using the test operators ‘-a’ and ‘-o’. On System V, the precedence of ‘-a’ and ‘-o’ is wrong relative to the unary operators; consequently, POSIX does not specify them, so using them is nonportable. If you combine ‘&&’ and ‘||’ in the same statement, keep in mind that they have equal precedence.

To enable configure scripts to support cross-compilation, they shouldn’t do anything that tests features of the host system instead of the target system. But occasionally you may find it necessary to check whether some arbitrary file exists. To do so, use ‘test -f’ or ‘test -r’. Do not use ‘test -x’, because 4.3BSD does not have it.

Another nonportable shell programming construction is

var=${var:-value} The intent is to set var to value only if it is not already set, but if var has any value, even the empty string, to leave it alone. Old BSD shells, including the Ultrix sh, don’t accept the colon, and complain and die. A portable equivalent is

: ${var=value}

5.7 Multiple Cases

Some operations are accomplished in several possible ways, depending on the UNIX variant. Checking for them essentially requires a “case statement”. Autoconf does not directly provide one; however, it is easy to simulate by using a shell variable to keep track of whether a way to perform the operation has been found yet.

Here is an example that uses the shell variable fstype to keep track of whether the remaining cases need to be checked.

AC_MSG_CHECKING(how to get filesystem type)
fstype=no
# The order of these tests is important.
AC_TRY_CPP([#include <sys/statvfs.h>
#include <sys/fstyp.h>], AC_DEFINE(FSTYPE_STATVFS) fstype=SVR4)
if test $fstype = no; then
AC_TRY_CPP([#include <sys/statfs.h>
#include <sys/fstyp.h>], AC_DEFINE(FSTYPE_USG_STATFS) fstype=SVR3)
fi
if test $fstype = no; then
AC_TRY_CPP([#include <sys/statfs.h>
#include <sys/vmount.h>], AC_DEFINE(FSTYPE_AIX_STATFS) fstype=AIX)
fi
# (more cases omitted here)
AC_MSG_RESULT($fstype)

5.8 Language Choice

Packages that use both C and C++ need to test features of both compilers. Autoconfgenerated configure scripts check for C features by default. The following macros determine which language’s compiler is used in tests that follow in ‘configure.in’.

Macro AC LANG C

Do compilation tests using CC and CPP and use extension ‘.c’ for test programs. Set the shell variable cross_compiling to the value computed by AC_PROG_CC if it has been run, empty otherwise.

Macro AC LANG CPLUSPLUS

Do compilation tests using CXX and CXXCPP and use extension ‘.C’ for test programs. Set the shell variable cross_compiling to the value computed by AC_PROG_CXX if it has been run, empty otherwise.

Macro AC LANG FORTRAN77

Do compilation tests using F77 and use extension ‘.f’ for test programs. Set the shell variable cross_compiling to the value computed by AC_PROG_F77 if it has been run, empty otherwise.

Macro AC LANG SAVE

Remember the current language (as set by AC_LANG_C, AC_LANG_CPLUSPLUS or AC_LANG_FORTRAN77) on a stack. Does not change which language is current. Use this macro and AC_LANG_RESTORE in macros that need to temporarily switch to a particular language.

Macro AC LANG RESTORE

Select the language that is saved on the top of the stack, as set by AC_LANG_ SAVE, and remove it from the stack. This macro is equivalent to either AC_

LANG_C, AC_LANG_CPLUSPLUS or AC_LANG_FORTRAN77, whichever had been run

most recently when AC_LANG_SAVE was last called. Do not call this macro more times than AC_LANG_SAVE.

Macro AC REQUIRE CPP

Ensure that whichever preprocessor would currently be used for tests has been found. Calls AC_REQUIRE (see Section 7.4.1 [Prerequisite Macros], page 54) with an argument of either AC_PROG_CPP or AC_PROG_CXXCPP, depending on which language is current.

6 Results of Tests

Once configure has determined whether a feature exists, what can it do to record that information? There are four sorts of things it can do: define a C preprocessor symbol, set a variable in the output files, save the result in a cache file for future configure runs, and print a message letting the user know the result of the test.

6.1 Defining C Preprocessor Symbols

A common action to take in response to a feature test is to define a C preprocessor symbol indicating the results of the test. That is done by calling AC_DEFINE or AC_DEFINE_

UNQUOTED.

By default, AC_OUTPUT places the symbols defined by these macros into the output variable DEFS, which contains an option ‘-Dsymbol=value’ for each symbol defined. Unlike in Autoconf version 1, there is no variable DEFS defined while configure is running. To check whether Autoconf macros have already defined a certain C preprocessor symbol, test the value of the appropriate cache variable, as in this example:

AC_CHECK_FUNC(vprintf, AC_DEFINE(HAVE_VPRINTF))
if test "$ac_cv_func_vprintf" != yes; then
AC_CHECK_FUNC(_doprnt, AC_DEFINE(HAVE_DOPRNT))
fi

If AC_CONFIG_HEADER has been called, then instead of creating DEFS, AC_OUTPUT creates a header file by substituting the correct values into #define statements in a template file. See Section 3.4 [Configuration Headers], page 15, for more information about this kind of output.

Macro AC DEFINE (variable [, value [, description]])

Define C preprocessor variable variable. If value is given, set variable to that value (verbatim), otherwise set it to 1. value should not contain literal newlines, and if you are not using AC_CONFIG_HEADER it should not contain any ‘#’ characters, as make tends to eat them. To use a shell variable (which you need to do in order to define a value containing the m4 quote characters ‘[’ or ‘]’), use AC_DEFINE_UNQUOTED instead. description is only useful if you are using AC_CONFIG_HEADER. In this case, description is put into the generated ‘config.h.in’ as the comment before the macro define; the macro need not be mentioned in ‘acconfig.h’. The following example defines the C preprocessor variable EQUATION to be the string constant ‘"$a > $b"’:

     AC_DEFINE(EQUATION, "$a > $b")

Macro AC DEFINE UNQUOTED (variable [, value [, description]])

Like AC_DEFINE, but three shell expansions are performed—once—on variable and value: variable expansion (‘$’), command substitution (‘`’), and backslash escaping (‘\’). Single and double quote characters in the value have no special meaning. Use this macro instead of AC_DEFINE when variable or value is a shell variable. Examples:

     AC_DEFINE_UNQUOTED(config_machfile, "${machfile}")
     AC_DEFINE_UNQUOTED(GETGROUPS_T, $ac_cv_type_getgroups)
     AC_DEFINE_UNQUOTED(${ac_tr_hdr})

Due to the syntactical bizarreness of the Bourne shell, do not use semicolons to separate AC_DEFINE or AC_DEFINE_UNQUOTED calls from other macro calls or shell code; that can cause syntax errors in the resulting configure script. Use either spaces or newlines. That is, do this:

AC_CHECK_HEADER(elf.h, AC_DEFINE(SVR4) LIBS="$LIBS -lelf")

or this:

AC_CHECK_HEADER(elf.h,
  AC_DEFINE(SVR4)
  LIBS="$LIBS -lelf")

instead of this:

AC_CHECK_HEADER(elf.h, AC_DEFINE(SVR4); LIBS="$LIBS -lelf")

6.2 Setting Output Variables

One way to record the results of tests is to set output variables, which are shell variables whose values are substituted into files that configure outputs. The two macros below create new output variables. See Section 3.3.1 [Preset Output Variables], page 11, for a list of output variables that are always available.

Macro AC SUBST (variable)

Create an output variable from a shell variable. Make AC_OUTPUT substitute the variable variable into output files (typically one or more ‘Makefile’s). This means that AC_OUTPUT will replace instances of ‘@variable@’ in input files with the value that the shell variable variable has when AC_OUTPUT is called. The value of variable should not contain literal newlines.

Macro AC SUBST FILE (variable)

Another way to create an output variable from a shell variable. Make AC_OUTPUT insert (without substitutions) the contents of the file named by shell variable variable into output files. This means that AC_OUTPUT will replace instances of ‘@variable@’ in output files (such as ‘Makefile.in’) with the contents of the file that the shell variable variable names when AC_OUTPUT is called. Set the variable to ‘/dev/null’ for cases that do not have a file to insert.

This macro is useful for inserting ‘Makefile’ fragments containing special dependencies or other make directives for particular host or target types into ‘Makefile’s. For example, ‘configure.in’ could contain:

     AC_SUBST_FILE(host_frag)dnl
     host_frag=$srcdir/conf/sun4.mh

and then a ‘Makefile.in’ could contain:

     @host_frag@

6.3 Caching Results

To avoid checking for the same features repeatedly in various configure scripts (or repeated runs of one script), configure saves the results of many of its checks in a cache file. If, when a configure script runs, it finds a cache file, it reads from it the results from previous runs and avoids rerunning those checks. As a result, configure can run much faster than if it had to perform all of the checks every time.

Macro AC CACHE VAL (cache-id, commands-to-set-it)

Ensure that the results of the check identified by cache-id are available. If the results of the check were in the cache file that was read, and configure was not given the ‘--quiet’ or ‘--silent’ option, print a message saying that the result was cached; otherwise, run the shell commands commands-to-set-it. Those commands should have no side effects except for setting the variable cache-id. In particular, they should not call AC_DEFINE; the code that follows the call to AC_CACHE_VAL should do that, based on the cached value. Also, they should not print any messages, for example with AC_MSG_CHECKING; do that before calling AC_CACHE_VAL, so the messages are printed regardless of whether the results of the check are retrieved from the cache or determined by running the shell commands. If the shell commands are run to determine the value, the value will be saved in the cache file just before configure creates its output files. See Section 6.3.1 [Cache Variable Names], page 48, for how to choose the name of the cache-id variable.

Macro AC CACHE CHECK (message, cache-id, commands)

A wrapper for AC_CACHE_VAL that takes care of printing the messages. This macro provides a convenient shorthand for the most common way to use these macros. It calls AC_MSG_CHECKING for message, then AC_CACHE_VAL with the cache-id and commands arguments, and AC_MSG_RESULT with cache-id.

Macro AC CACHE LOAD

Loads values from existing cache file, or creates a new cache file if a cache file is not found. Called automatically from AC_INIT.

Macro AC CACHE SAVE

Flushes all cached values to the cache file. Called automatically from AC_ OUTPUT, but it can be quite useful to call AC_CACHE_SAVE at key points in configure.in. Doing so checkpoints the cache in case of an early configure script abort.

6.3.1 Cache Variable Names

The names of cache variables should have the following format:

package-prefix_cv_value-type_specific-value[_additional-options]

for example, ‘ac_cv_header_stat_broken’ or ‘ac_cv_prog_gcc_traditional’. The parts

of the variable name are: package-prefix

An abbreviation for your package or organization; the same prefix you begin local Autoconf macros with, except lowercase by convention. For cache values used by the distributed Autoconf macros, this value is ‘ac’.

_cv_Indicates that this shell variable is a cache value.

value-type A convention for classifying cache values, to produce a rational naming system. The values used in Autoconf are listed in Section 7.2 [Macro Names], page 52.

specific-value Which member of the class of cache values this test applies to. For example, which function (‘alloca’), program (‘gcc’), or output variable (‘INSTALL’).

additional-options Any particular behavior of the specific member that this test applies to. For example, ‘broken’ or ‘set’. This part of the name may be omitted if it does not apply.

The values assigned to cache variables may not contain newlines. Usually, their values will be boolean (‘yes’ or ‘no’) or the names of files or functions; so this is not an important restriction.

6.3.2 Cache Files

A cache file is a shell script that caches the results of configure tests run on one system so they can be shared between configure scripts and configure runs. It is not useful on other systems. If its contents are invalid for some reason, the user may delete or edit it.

By default, configure uses ‘./config.cache’ as the cache file, creating it if it does not exist already. configure accepts the ‘--cache-file=file’ option to use a different cache file; that is what configure does when it calls configure scripts in subdirectories, so they share the cache. See Section 3.5 [Subdirectories], page 17, for information on configuring subdirectories with the AC_CONFIG_SUBDIRS macro.

   Giving ‘--cache-file=/dev/null’ disables caching, for debugging configure. ‘config.status’

only pays attention to the cache file if it is given the ‘--recheck’ option, which makes it rerun configure. If you are anticipating a long debugging period, you can also disable cache loading and saving for a configure script by redefining the cache macros at the start

of ‘configure.in’:
     define([AC_CACHE_LOAD], )dnl
     define([AC_CACHE_SAVE], )dnl
     AC_INIT(whatever)

... rest of configure.in ... It is wrong to try to distribute cache files for particular system types. There is too

much room for error in doing that, and too much administrative overhead in maintaining them. For any features that can’t be guessed automatically, use the standard method of the canonical system type and linking files (see Chapter 8 [Manual Configuration], page 56).

The cache file on a particular system will gradually accumulate whenever someone runs a configure script; it will be initially nonexistent. Running configure merges the new cache results with the existing cache file. The site initialization script can specify a site-wide cache file to use instead of the default, to make it work transparently, as long as the same C compiler is used every time (see Section 9.5 [Site Defaults], page 63).

If your configure script, or a macro called from configure.in, happens to abort the configure process, it may be useful to checkpoint the cache a few times at key points. Doing so will reduce the amount of time it takes to re-run the configure script with (hopefully) the error that caused the previous abort corrected.

... AC INIT, etc. ...

dnl checks for programs
AC_PROG_CC
AC_PROG_GCC_TRADITIONAL

... more program checks ...

AC_CACHE_SAVE
dnl checks for libraries
AC_CHECK_LIB(nsl, gethostbyname)
AC_CHECK_LIB(socket, connect)

... more lib checks ...

AC_CACHE_SAVE
dnl Might abort...
AM_PATH_GTK(1.0.2, , exit 1)
AM_PATH_GTKMM(0.9.5, , exit 1)

6.4 Printing Messages

configure scripts need to give users running them several kinds of information. The following macros print messages in ways appropriate for each kind. The arguments to all of them get enclosed in shell double quotes, so the shell performs variable and backquote substitution on them. You can print a message containing a comma by quoting the message with the m4 quote characters:

AC_MSG_RESULT([never mind, I found the BASIC compiler])

These macros are all wrappers around the echo shell command. configure scripts should rarely need to run echo directly to print messages for the user. Using these macros makes it easy to change how and when each kind of message is printed; such changes need only be made to the macro definitions, and all of the callers change automatically.

Macro AC MSG CHECKING (feature-description)

Notify the user that configure is checking for a particular feature. This macro prints a message that starts with ‘checking ’ and ends with ‘...’ and no newline. It must be followed by a call to AC_MSG_RESULT to print the result of the check and the newline. The feature-description should be something like

‘whether the Fortran compiler accepts C++ comments’ or ‘for c89’.

This macro prints nothing if configure is run with the ‘--quiet’ or ‘--silent’ option.

Macro AC MSG RESULT (result-description)

Notify the user of the results of a check. result-description is almost always the value of the cache variable for the check, typically ‘yes’, ‘no’, or a file name. This macro should follow a call to AC_MSG_CHECKING, and the resultdescription should be the completion of the message printed by the call to

AC_MSG_CHECKING.

This macro prints nothing if configure is run with the ‘--quiet’ or ‘--silent’ option.

Macro AC MSG ERROR (error-description)

Notify the user of an error that prevents configure from completing. This macro prints an error message on the standard error output and exits configure with a nonzero status. error-description should be something like ‘invalid

value $HOME for \$HOME’.

Macro AC MSG WARN (problem-description)

Notify the configure user of a possible problem. This macro prints the message on the standard error output; configure continues running afterward, so macros that call AC_MSG_WARN should provide a default (back-up) behavior for the situations they warn about. problem-description should be something like

‘ln -s seems to make hard links’.

The following two macros are an obsolete alternative to AC_MSG_CHECKING and AC_MSG_

RESULT.

Macro AC CHECKING (feature-description)

This macro is similar to AC_MSG_CHECKING, except that it prints a newline after the feature-description. It is useful mainly to print a general description of the overall purpose of a group of feature checks, e.g.,

     AC_CHECKING(if stack overflow is detectable)

Macro AC VERBOSE (result-description)

This macro is similar to AC_MSG_RESULT, except that it is meant to follow a call to AC_CHECKING instead of AC_MSG_CHECKING; it starts the message it prints with a tab. It is considered obsolete.

7 Writing Macros

When you write a feature test that could be applicable to more than one software package, the best thing to do is encapsulate it in a new macro. Here are some instructions and guidelines for writing Autoconf macros.

7.1 Macro Definitions

Autoconf macros are defined using the AC_DEFUN macro, which is similar to the m4 builtin define macro. In addition to defining a macro, AC_DEFUN adds to it some code which is used to constrain the order in which macros are called (see Section 7.4.1 [Prerequisite Macros], page 54).

An Autoconf macro definition looks like this:

AC_DEFUN(macro-name, [macro-body]) The square brackets here do not indicate optional text: they should literally be present in the macro definition to avoid macro expansion problems (see Section 7.3 [Quoting], page 53). You can refer to any arguments passed to the macro as ‘$1’, ‘$2’, etc.

To introduce comments in m4, use the m4 builtin dnl; it causes m4 to discard the text through the next newline. It is not needed between macro definitions in ‘acsite.m4’ and ‘aclocal.m4’, because all output is discarded until AC_INIT is called.

See section “How to define new macros” in GNU m4, for more complete information on writing m4 macros.

7.2 Macro Names

All of the Autoconf macros have all-uppercase names starting with ‘AC_’ to prevent them from accidentally conflicting with other text. All shell variables that they use for internal purposes have mostly-lowercase names starting with ‘ac_’. To ensure that your macros don’t conflict with present or future Autoconf macros, you should prefix your own macro names and any shell variables they use with some other sequence. Possibilities include your initials, or an abbreviation for the name of your organization or software package.

Most of the Autoconf macros’ names follow a structured naming convention that indicates the kind of feature check by the name. The macro names consist of several words, separated by underscores, going from most general to most specific. The names of their cache variables use the same convention (see Section 6.3.1 [Cache Variable Names], page 48, for more information on them).

The first word of the name after ‘AC_’ usually tells the category of feature being tested. Here are the categories used in Autoconf for specific test macros, the kind of macro that you are more likely to write. They are also used for cache variables, in all-lowercase. Use them where applicable; where they’re not, invent your own categories.

C C language builtin features.

DECL Declarations of C variables in header files.

FUNC Functions in libraries.

GROUP UNIX group owners of files.

HEADER Header files.

LIB C libraries.

PATH The full path names to files, including programs.

PROG The base names of programs.

STRUCT Definitions of C structures in header files.

SYS Operating system features.

TYPE C builtin or declared types.

VAR C variables in libraries.

After the category comes the name of the particular feature being tested. Any further words in the macro name indicate particular aspects of the feature. For example, AC_FUNC_ UTIME_NULL checks the behavior of the utime function when called with a NULL pointer.

A macro that is an internal subroutine of another macro should have a name that starts with the name of that other macro, followed by one or more words saying what the internal macro does. For example, AC_PATH_X has internal macros AC_PATH_X_XMKMF and AC_PATH_

X_DIRECT.

7.3 Quoting

Macros that are called by other macros are evaluated by m4 several times; each evaluation might require another layer of quotes to prevent unwanted expansions of macros or m4 builtins, such as ‘define’ and ‘$1’. Quotes are also required around macro arguments that contain commas, since commas separate the arguments from each other. It’s a good idea to quote any macro arguments that contain newlines or calls to other macros, as well.

Autoconf changes the m4 quote characters from the default ‘`’ and ‘'’ to ‘[’ and ‘]’, because many of the macros use ‘`’ and ‘'’, mismatched. However, in a few places the macros need to use brackets (usually in C program text or regular expressions). In those places, they use the m4 builtin command changequote to temporarily change the quote characters to ‘<<’ and ‘>>’. (Sometimes, if they don’t need to quote anything, they disable quoting entirely instead by setting the quote characters to empty strings.) Here is an example:

     AC_TRY_LINK(
     changequote(<<, >>)dnl
     <<#include <time.h>
     #ifndef tzname /* For SGI.  */
     extern char *tzname[]; /* RS6000 and others reject char **tzname.  */
     #endif>>,
     changequote([, ])dnl
     [atoi(*tzname);], ac_cv_var_tzname=yes, ac_cv_var_tzname=no)

When you create a configure script using newly written macros, examine it carefully to check whether you need to add more quotes in your macros. If one or more words have disappeared in the m4 output, you need more quotes. When in doubt, quote.

However, it’s also possible to put on too many layers of quotes. If this happens, the resulting configure script will contain unexpanded macros. The autoconf program checks for this problem by doing ‘grep AC_configure’.

7.4 Dependencies Between Macros

Some Autoconf macros depend on other macros having been called first in order to work correctly. Autoconf provides a way to ensure that certain macros are called if needed and a way to warn the user if macros are called in an order that might cause incorrect operation.

7.4.1 Prerequisite Macros

A macro that you write might need to use values that have previously been computed by other macros. For example, AC_DECL_YYTEXT examines the output of flex or lex, so it depends on AC_PROG_LEX having been called first to set the shell variable LEX.

Rather than forcing the user of the macros to keep track of the dependencies between them, you can use the AC_REQUIRE macro to do it automatically. AC_REQUIRE can ensure that a macro is only called if it is needed, and only called once.

Macro AC REQUIRE (macro-name)

If the m4 macro macro-name has not already been called, call it (without any arguments). Make sure to quote macro-name with square brackets. macroname must have been defined using AC_DEFUN or else contain a call to AC_ PROVIDE to indicate that it has been called.

An alternative to using AC_DEFUN is to use define and call AC_PROVIDE. Because this technique does not prevent nested messages, it is considered obsolete.

Macro AC PROVIDE (this-macro-name)

Record the fact that this-macro-name has been called. this-macro-name should be the name of the macro that is calling AC_PROVIDE. An easy way to get it is from the m4 builtin variable $0, like this:

          AC_PROVIDE([$0])

7.4.2 Suggested Ordering

Some macros should be run before another macro if both are called, but neither requires that the other be called. For example, a macro that changes the behavior of the C compiler should be called before any macros that run the C compiler. Many of these dependencies are noted in the documentation.

Autoconf provides the AC_BEFORE macro to warn users when macros with this kind of dependency appear out of order in a ‘configure.in’ file. The warning occurs when creating

configure from ‘configure.in’, not when running configure. For example, AC_PROG_CPP

checks whether the C compiler can run the C preprocessor when given the ‘-E’ option. It should therefore be called after any macros that change which C compiler is being used,

such as AC_PROG_CC. So AC_PROG_CC contains:

AC_BEFORE([$0], [AC_PROG_CPP])dnl

This warns the user if a call to AC_PROG_CPP has already occurred when AC_PROG_CC is called.

Macro AC BEFORE (this-macro-name, called-macro-name)

Make m4 print a warning message on the standard error output if called-macro-name has already been called. this-macro-name should be the name of the macro that is calling AC_BEFORE. The macro called-macro-name must have been defined using AC_DEFUN or else contain a call to AC_PROVIDE to indicate that it has been called.

7.4.3 Obsolete Macros

Configuration and portability technology has evolved over the years. Often better ways of solving a particular problem are developed, or ad-hoc approaches are systematized. This process has occurred in many parts of Autoconf. One result is that some of the macros are now considered obsolete; they still work, but are no longer considered the best thing to do. Autoconf provides the AC_OBSOLETE macro to warn users producing configure scripts when they use obsolete macros, to encourage them to modernize. A sample call is:

AC_OBSOLETE([$0], [; use AC_CHECK_HEADERS(unistd.h) instead])dnl

Macro AC OBSOLETE (this-macro-name [, suggestion])

Make m4 print a message on the standard error output warning that this-macro-name is obsolete, and giving the file and line number where it was called. thismacro-name should be the name of the macro that is calling AC_OBSOLETE. If suggestion is given, it is printed at the end of the warning message; for example, it can be a suggestion for what to use instead of this-macro-name.

8 Manual Configuration

A few kinds of features can’t be guessed automatically by running test programs. For example, the details of the object file format, or special options that need to be passed to the compiler or linker. You can check for such features using ad-hoc means, such as having configure check the output of the uname program, or looking for libraries that are unique to particular systems. However, Autoconf provides a uniform method for handling unguessable features.

8.1 Specifying the System Type

Like other GNU configure scripts, Autoconf-generated configure scripts can make decisions based on a canonical name for the system type, which has the form:

cpu-company-system configure can usually guess the canonical name for the type of system it’s running on.

To do so it runs a script called config.guess, which derives the name using the uname command or symbols predefined by the C preprocessor.

Alternately, the user can specify the system type with command line arguments to configure. Doing so is necessary when cross-compiling. In the most complex case of cross-compiling, three system types are involved. The options to specify them are:

--build=build-type the type of system on which the package is being configured and compiled (rarely needed);

--host=host-type the type of system on which the package will run;

--target=target-type the type of system for which any compiler tools in the package will produce code.

If the user gives configure a non-option argument, it is used as the default for the host, target, and build system types if the user does not specify them explicitly with options. The target and build types default to the host type if it is given and they are not. If you are cross-compiling, you still have to specify the names of the cross-tools you use, in particular the C compiler, on the configure command line, e.g.,

     CC=m68k-coff-gcc configure --target=m68k-coff

configure recognizes short aliases for many system types; for example, ‘decstation’ can be given on the command line instead of ‘mips-dec-ultrix4.2’. configure runs a script called config.sub to canonicalize system type aliases.

8.2 Getting the Canonical System Type

The following macros make the system type available to configure scripts. They run the shell script config.guess to determine any values for the host, target, and build types that they need and the user did not specify on the command line. They run config.sub to canonicalize any aliases the user gave. If you use these macros, you must distribute those two shell scripts along with your source code. See Section 3.2 [Output], page 9, for information about the AC_CONFIG_AUX_DIR macro which you can use to control which directory configure looks for those scripts in. If you do not use either of these macros, configure ignores any ‘--host’, ‘--target’, and ‘--build’ options given to it.

Macro AC CANONICAL SYSTEM

Determine the system type and set output variables to the names of the canonical system types. See Section 8.3 [System Type Variables], page 57, for details about the variables this macro sets.

Macro AC CANONICAL HOST

Perform only the subset of AC_CANONICAL_SYSTEM relevant to the host type. This is all that is needed for programs that are not part of a compiler toolchain.

Macro AC VALIDATE CACHED SYSTEM TUPLE (cmd)

If the cache file is inconsistent with the current host, target and build system types, execute cmd or print a default error message.

8.3 System Type Variables

After calling AC_CANONICAL_SYSTEM, the following output variables contain the system type information. After AC_CANONICAL_HOST, only the host variables below are set.

build, host, target

the canonical system names;

build_alias, host_alias, target_alias

the names the user specified, or the canonical names if config.guess was used;

build_cpu, build_vendor, build_os
host_cpu, host_vendor, host_os
target_cpu, target_vendor, target_os

the individual parts of the canonical names (for convenience).

8.4 Using the System Type

How do you use a canonical system type? Usually, you use it in one or more case statements in ‘configure.in’ to select system-specific C files. Then link those files, which have names based on the system name, to generic names, such as ‘host.h’ or ‘target.c’. The case statement patterns can use shell wildcards to group several cases together, like in this fragment:

     case "$target" in
     i386-*-mach* | i386-*-gnu*) obj_format=aout emulation=mach bfd_gas=yes ;;
     i960-*-bout) obj_format=bout ;;
     esac

Macro AC LINK FILES (source…, dest…)

Make AC_OUTPUT link each of the existing files source to the corresponding link name dest. Makes a symbolic link if possible, otherwise a hard link. The dest and source names should be relative to the top level source or build directory. This macro may be called multiple times. For example, this call:

     AC_LINK_FILES(config/${machine}.h config/${obj_format}.h, host.h object.h)

creates in the current directory ‘host.h’, which is a link to ‘srcdir/config/${machine}.h’,

and ‘object.h’, which is a link to ‘srcdir/config/${obj_format}.h’.

You can also use the host system type to find cross-compilation tools. See Section 4.1.2 [Generic Programs], page 22, for information about the AC_CHECK_TOOL macro which does that.

9 Site Configuration

configure scripts support several kinds of local configuration decisions. There are ways for users to specify where external software packages are, include or exclude optional features, install programs under modified names, and set default values for configure options.

9.1 Working With External Software

Some packages require, or can optionally use, other software packages which are already installed. The user can give configure command line options to specify which such external software to use. The options have one of these forms:

--with-package[=arg]

     --without-package

For example, ‘--with-gnu-ld’ means work with the GNU linker instead of some other linker. ‘--with-x’ means work with The X Window System.

The user can give an argument by following the package name with ‘=’ and the argument. Giving an argument of ‘no’ is for packages that are used by default; it says to not use the package. An argument that is neither ‘yes’ nor ‘no’ could include a name or number of a version of the other package, to specify more precisely which other package this program is supposed to work with. If no argument is given, it defaults to ‘yes’. ‘--without-package’ is equivalent to ‘--with-package=no’.

configure scripts do not complain about ‘--with-package’ options that they do not support. This behavior permits configuring a source tree containing multiple packages with a top-level configure script when the packages support different options, without spurious error messages about options that some of the packages support. An unfortunate side effect is that option spelling errors are not diagnosed. No better approach to this problem has been suggested so far.

For each external software package that may be used, ‘configure.in’ should call AC_ ARG_WITH to detect whether the configure user asked to use it. Whether each package is used or not by default, and which arguments are valid, is up to you.

Macro AC ARG WITH (package, help-string [, action-if-given [,

action-if-not-given]]) If the user gave configure the option ‘--with-package’ or ‘--without-package’, run shell commands action-if-given. If neither option was given, run shell commands action-if-not-given. The name package indicates another software package that this program should work with. It should consist only of alphanumeric characters and dashes. The option’s argument is available to the shell commands action-if-given in the shell variable withval, which is actually just the value of the shell variable with_package, with any ‘-’ characters changed into ‘_’. You may use that variable instead, if you wish. The argument help-string is a description of the option which looks like this:

            --with-readline         support fancy command line editing

help-string may be more than one line long, if more detail is needed. Just make sure the columns line up in ‘configure --help’. Avoid tabs in the help string. You’ll need to enclose it in ‘[’ and ‘]’ in order to produce the leading spaces.

Macro AC WITH (package, action-if-given [, action-if-not-given])

This is an obsolete version of AC_ARG_WITH that does not support providing a help string.

9.2 Choosing Package Options

If a software package has optional compile-time features, the user can give configure command line options to specify whether to compile them. The options have one of these forms:

--enable-feature[=arg]

     --disable-feature

These options allow users to choose which optional features to build and install. ‘--enable-feature’ options should never make a feature behave differently or cause one feature to replace another. They should only cause parts of the program to be built rather than left out.

The user can give an argument by following the feature name with ‘=’ and the argument. Giving an argument of ‘no’ requests that the feature not be made available. A feature with an argument looks like ‘--enable-debug=stabs’. If no argument is given, it defaults to ‘yes’. ‘--disable-feature’ is equivalent to ‘--enable-feature=no’.

configure scripts do not complain about ‘--enable-feature’ options that they do not support. This behavior permits configuring a source tree containing multiple packages with a top-level configure script when the packages support different options, without spurious error messages about options that some of the packages support. An unfortunate side effect is that option spelling errors are not diagnosed. No better approach to this problem has been suggested so far.

For each optional feature, ‘configure.in’ should call AC_ARG_ENABLE to detect whether the configure user asked to include it. Whether each feature is included or not by default, and which arguments are valid, is up to you.

Macro AC ARG ENABLE (feature, help-string [, action-if-given [,

action-if-not-given]]) If the user gave configure the option ‘--enable-feature’ or ‘--disable-feature’, run shell commands action-if-given. If neither option was given, run shell commands action-if-not-given. The name feature indicates an optional user-level facility. It should consist only of alphanumeric characters and dashes. The option’s argument is available to the shell commands action-if-given in the shell variable enableval, which is actually just the value of the shell variable enable_feature, with any ‘-’ characters changed into ‘_’. You may use that variable instead, if you wish. The help-string argument is like that of AC_ARG_ WITH (see Section 9.1 [External Software], page 59).

Macro AC ENABLE (feature, action-if-given [, action-if-not-given])

This is an obsolete version of AC_ARG_ENABLE that does not support providing a help string.

9.3 Configuring Site Details

Some software packages require complex site-specific information. Some examples are host names to use for certain services, company names, and email addresses to contact. Since some configuration scripts generated by Metaconfig ask for such information interactively, people sometimes wonder how to get that information in Autoconf-generated configuration scripts, which aren’t interactive.

Such site configuration information should be put in a file that is edited only by users, not by programs. The location of the file can either be based on the prefix variable, or be a standard location such as the user’s home directory. It could even be specified by an environment variable. The programs should examine that file at run time, rather than at compile time. Run time configuration is more convenient for users and makes the configuration process simpler than getting the information while configuring. See section “Variables for Installation Directories” in GNU Coding Standards, for more information on where to put data files.

9.4 Transforming Program Names When Installing

Autoconf supports changing the names of programs when installing them. In order to use these transformations, ‘configure.in’ must call the macro AC_ARG_PROGRAM.

Macro AC ARG PROGRAM

Place in output variable program_transform_name a sequence of sed commands for changing the names of installed programs.

If any of the options described below are given to configure, program names are transformed accordingly. Otherwise, if AC_CANONICAL_SYSTEM has been called and a ‘--target’ value is given that differs from the host type (specified with ‘--host’ or defaulted by config.sub), the target type followed by a dash is used as a prefix. Otherwise, no program name transformation is done.

9.4.1 Transformation Options

You can specify name transformations by giving configure these command line options:

--program-prefix=prefix

prepend prefix to the names;

--program-suffix=suffix

append suffix to the names;

--program-transform-name=expression

perform sed substitution expression on the names.

9.4.2 Transformation Examples

These transformations are useful with programs that can be part of a cross-compilation development environment. For example, a cross-assembler running on a Sun 4 configured with ‘--target=i960-vxworks’ is normally installed as ‘i960-vxworks-as’, rather than ‘as’, which could be confused with a native Sun 4 assembler.

You can force a program name to begin with ‘g’, if you don’t want GNU programs installed on your system to shadow other programs with the same name. For example, if you configure GNU diff with ‘--program-prefix=g’, then when you run ‘make install’

it is installed as ‘/usr/local/bin/gdiff’.

As a more sophisticated example, you could use

     --program-transform-name='s/^/g/; s/^gg/g/; s/^gless/less/'

to prepend ‘g’ to most of the program names in a source tree, excepting those like gdb that already have one and those like less and lesskey that aren’t GNU programs. (That is assuming that you have a source tree containing those programs that is set up to use this feature.)

One way to install multiple versions of some programs simultaneously is to append a version number to the name of one or both. For example, if you want to keep Autoconf version 1 around for awhile, you can configure Autoconf version 2 using ‘--program-suffix=2’

to install the programs as ‘/usr/local/bin/autoconf2’, ‘/usr/local/bin/autoheader2’,

etc.

9.4.3 Transformation Rules

   Here is how to use the variable program_transform_name in a ‘Makefile.in’:
     transform=@program_transform_name@
     install: all
             $(INSTALL_PROGRAM) myprog $(bindir)/`echo myprog|sed '$(transform)'`
     uninstall:
             rm -f $(bindir)/`echo myprog|sed '$(transform)'`

If you have more than one program to install, you can do it in a loop:

     PROGRAMS=cp ls rm
     install:
             for p in $(PROGRAMS); do \
               $(INSTALL_PROGRAM) $$p $(bindir)/`echo $$p|sed '$(transform)'`; \
             done
     uninstall:
             for p in $(PROGRAMS); do \
               rm -f $(bindir)/`echo $$p|sed '$(transform)'`; \
             done

Whether to do the transformations on documentation files (Texinfo or man) is a tricky question; there seems to be no perfect answer, due to the several reasons for name transforming. Documentation is not usually particular to a specific architecture, and Texinfo files do not conflict with system documentation. But they might conflict with earlier versions of the same files, and man pages sometimes do conflict with system documentation. As a compromise, it is probably best to do name transformations on man pages but not on Texinfo manuals.

9.5 Setting Site Defaults

Autoconf-generated configure scripts allow your site to provide default values for some configuration values. You do this by creating site- and system-wide initialization files.

If the environment variable CONFIG_SITE is set, configure uses its value as the name of a shell script to read. Otherwise, it reads the shell script ‘prefix/share/config.site’ if it exists, then ‘prefix/etc/config.site’ if it exists. Thus, settings in machine-specific files override those in machine-independent ones in case of conflict.

Site files can be arbitrary shell scripts, but only certain kinds of code are really appropriate to be in them. Because configure reads any cache file after it has read any site files, a site file can define a default cache file to be shared between all Autoconf-generated configure scripts run on that system. If you set a default cache file in a site file, it is a good idea to also set the output variable CC in that site file, because the cache file is only valid for a particular compiler, but many systems have several available.

You can examine or override the value set by a command line option to configure in a site file; options set shell variables that have the same names as the options, with any dashes turned into underscores. The exceptions are that ‘--without-’ and ‘--disable-’ options are like giving the corresponding ‘--with-’ or ‘--enable-’ option and the value ‘no’.

Thus, ‘--cache-file=localcache’ sets the variable cache_file to the value ‘localcache’;
‘--enable-warnings=no’ or ‘--disable-warnings’ sets the variable enable_warnings to

the value ‘no’; ‘--prefix=/usr’ sets the variable prefix to the value ‘/usr’; etc.

Site files are also good places to set default values for other output variables, such as CFLAGS, if you need to give them non-default values: anything you would normally do, repetitively, on the command line. If you use non-default values for prefix or exec prefix (wherever you locate the site file), you can set them in the site file if you specify it with the CONFIG_SITE environment variable.

You can set some cache values in the site file itself. Doing this is useful if you are cross-compiling, so it is impossible to check features that require running a test program. You could “prime the cache” by setting those values correctly for that system in ‘prefix/etc/config.site’. To find out the names of the cache variables you need to set, look for shell variables with ‘_cv_’ in their names in the affected configure scripts, or in the Autoconf m4 source code for those macros.

The cache file is careful to not override any variables set in the site files. Similarly, you should not override command-line options in the site files. Your code should check that variables such as prefix and cache_file have their default values (as set near the top of configure) before changing them.

   Here is a sample file ‘/usr/share/local/gnu/share/config.site’. The command
‘configure --prefix=/usr/share/local/gnu’ would read this file (if CONFIG_SITE is not

set to a different file).

# config.site for configure
#
# Change some defaults.
test "$prefix" = NONE && prefix=/usr/share/local/gnu
test "$exec_prefix" = NONE && exec_prefix=/usr/local/gnu
test "$sharedstatedir" = '${prefix}/com' && sharedstatedir=/var
test "$localstatedir" = '${prefix}/var' && localstatedir=/var
#
# Give Autoconf 2.x generated configure scripts a shared default
# cache file for feature test results, architecture-specific.
if test "$cache_file" = ./config.cache; then
  cache_file="$prefix/var/config.cache"
  # A cache file is only valid for one C compiler.
  CC=gcc
fi

10 Running configure Scripts

Below are instructions on how to configure a package that uses a configure script, suitable for inclusion as an ‘INSTALL’ file in the package. A plain-text version of ‘INSTALL’ which you may use comes with Autoconf.

Invoking config.status, Questions, Invoking configure, Top

11 Recreating a Configuration

The configure script creates a file named ‘config.status’ which describes which configuration options were specified when the package was last configured. This file is a shell script which, if run, will recreate the same configuration.

You can give ‘config.status’ the ‘--recheck’ option to update itself. This option is useful if you change configure, so that the results of some tests might be different from the previous run. The ‘--recheck’ option re-runs configure with the same arguments you used before, plus the ‘--no-create’ option, which prevent configure from running ‘config.status’ and creating ‘Makefile’ and other files, and the ‘--no-recursion’ option, which prevents configure from running other configure scripts in subdirectories. (This is so other ‘Makefile’ rules can run ‘config.status’ when it changes; see Section 3.3.3 [Automatic Remaking], page 14, for an example).

‘config.status’ also accepts the options ‘--help’, which prints a summary of the options to ‘config.status’, and ‘--version’, which prints the version of Autoconf used to create the configure script that generated ‘config.status’.

‘config.status’ checks several optional environment variables that can alter its behavior:

Variable CONFIG SHELL

The shell with which to run configure for the ‘--recheck’ option. It must be Bourne-compatible. The default is ‘/bin/sh’.

Variable CONFIG STATUS

The file name to use for the shell script that records the configuration. The default is ‘./config.status’. This variable is useful when one package uses parts of another and the configure scripts shouldn’t be merged because they are maintained separately.

The following variables provide one way for separately distributed packages to share the values computed by configure. Doing so can be useful if some of the packages need a superset of the features that one of them, perhaps a common library, does. These variables allow a ‘config.status’ file to create files other than the ones that its ‘configure.in’ specifies, so it can be used for a different package.

Variable CONFIG FILES

The files in which to perform ‘@variable@’ substitutions. The default is the

     arguments given to AC_OUTPUT in ‘configure.in’.

Variable CONFIG HEADERS

The files in which to substitute C #define statements. The default is the arguments given to AC_CONFIG_HEADER; if that macro was not called, ‘config.status’ ignores this variable.

These variables also allow you to write ‘Makefile’ rules that regenerate only some of the files. For example, in the dependencies given above (see Section 3.3.3 [Automatic Remaking], page 14), ‘config.status’ is run twice when ‘configure.in’ has changed. If that bothers you, you can make each run only regenerate the files for that rule:

     config.h: stamp-h
     stamp-h: config.h.in config.status
             CONFIG_FILES= CONFIG_HEADERS=config.h ./config.status
             echo > stamp-h
     Makefile: Makefile.in config.status
             CONFIG_FILES=Makefile CONFIG_HEADERS= ./config.status
(If ‘configure.in’ does not call AC_CONFIG_HEADER, there is no need to set CONFIG_HEADERS

in the make rules.)

12 Questions About Autoconf

Several questions about Autoconf come up occasionally. Here some of them are addressed.

12.1 Distributing configure Scripts

What are the restrictions on distributing configure scripts that Autoconf generates? How does that affect my programs that use them?

There are no restrictions on how the configuration scripts that Autoconf produces may be distributed or used. In Autoconf version 1, they were covered by the GNU General Public License. We still encourage software authors to distribute their work under terms like those of the GPL, but doing so is not required to use Autoconf.

Of the other files that might be used with configure, ‘config.h.in’ is under whatever copyright you use for your ‘configure.in’, since it is derived from that file and from the public domain file ‘acconfig.h’. ‘config.sub’ and ‘config.guess’ have an exception to the GPL when they are used with an Autoconf-generated configure script, which permits you to distribute them under the same terms as the rest of your package. ‘install-sh’ is from the X Consortium and is not copyrighted.

12.2 Why Require GNU m4?

Why does Autoconf require GNU m4? Many m4 implementations have hard-coded limitations on the size and number of macros,

which Autoconf exceeds. They also lack several builtin macros that it would be difficult to get along without in a sophisticated application like Autoconf, including:

     builtin
     indir
     patsubst
     __file__
     __line__

Since only software maintainers need to use Autoconf, and since GNU m4 is simple to configure and install, it seems reasonable to require GNU m4 to be installed also. Many maintainers of GNU and other free software already have most of the GNU utilities installed, since they prefer them.

12.3 How Can I Bootstrap?

If Autoconf requires GNU m4 and GNU m4 has an Autoconf configure script, how do I bootstrap? It seems like a chicken and egg problem!

This is a misunderstanding. Although GNU m4 does come with a configure script produced by Autoconf, Autoconf is not required in order to run the script and install GNU m4. Autoconf is only required if you want to change the m4 configure script, which few people have to do (mainly its maintainer).

12.4 Why Not Imake?

Why not use Imake instead of configure scripts? Several people have written addressing this question, so I include adaptations of their

explanations here.

The following answer is based on one written by Richard Pixley:

Autoconf generated scripts frequently work on machines which it has never been set up to handle before. That is, it does a good job of inferring a configuration for a new system. Imake cannot do this.

Imake uses a common database of host specific data. For X11, this makes sense because the distribution is made as a collection of tools, by one central authority who has control over the database.

GNU tools are not released this way. Each GNU tool has a maintainer; these maintainers are scattered across the world. Using a common database would be a maintenance nightmare. Autoconf may appear to be this kind of database, but in fact it is not. Instead of listing host dependencies, it lists program requirements.

If you view the GNU suite as a collection of native tools, then the problems are similar. But the GNU development tools can be configured as cross tools in almost any host+target permutation. All of these configurations can be installed concurrently. They can even be configured to share host independent files across hosts. Imake doesn’t address these issues.

Imake templates are a form of standardization. The GNU coding standards address the same issues without necessarily imposing the same restrictions.

Here is some further explanation, written by Per Bothner:

One of the advantages of Imake is that it easy to generate large Makefiles using cpp’s ‘#include’ and macro mechanisms. However, cpp is not programmable: it has limited conditional facilities, and no looping. And cpp cannot inspect its environment.

All of these problems are solved by using sh instead of cpp. The shell is fully programmable, has macro substitution, can execute (or source) other shell scripts, and can inspect its environment.

Paul Eggert elaborates more:

With Autoconf, installers need not assume that Imake itself is already installed and working well. This may not seem like much of an advantage to people who are accustomed to Imake. But on many hosts Imake is not installed or the default installation is not working well, and requiring Imake to install a package hinders the acceptance of that package on those hosts. For example, the Imake template and configuration files might not be installed properly on a host, or the Imake build procedure might wrongly assume that all source files are in one big directory tree, or the Imake configuration might assume one compiler whereas the package or the installer needs to use another, or there might be a version mismatch between the Imake expected by the package and the Imake supported by the host. These problems are much rarer with Autoconf, where each package comes with its own independent configuration processor.

Also, Imake often suffers from unexpected interactions between make and the installer’s C preprocessor. The fundamental problem here is that the C preprocessor was designed to preprocess C programs, not ‘Makefile’s. This is much less of a problem with Autoconf, which uses the general-purpose preprocessor m4, and where the package’s author (rather than the installer) does the preprocessing in a standard way.

Finally, Mark Eichin notes:

Imake isn’t all that extensible, either. In order to add new features to Imake, you need to provide your own project template, and duplicate most of the features of the existing one. This means that for a sophisticated project, using the vendor-provided Imake templates fails to provide any leverage—since they don’t cover anything that your own project needs (unless it is an X11 program).

On the other side, though:

The one advantage that Imake has over configure: ‘Imakefile’s tend to be much shorter (likewise, less redundant) than ‘Makefile.in’s. There is a fix to this, however—at least for the Kerberos V5 tree, we’ve modified things to call in common ‘post.in’ and ‘pre.in’ ‘Makefile’ fragments for the entire tree. This means that a lot of common things don’t have to be duplicated, even though they normally are in configure setups.

13 Upgrading From Version 1

Autoconf version 2 is mostly backward compatible with version 1. However, it introduces better ways to do some things, and doesn’t support some of the ugly things in version 1. So, depending on how sophisticated your ‘configure.in’ files are, you might have to do some manual work in order to upgrade to version 2. This chapter points out some problems to watch for when upgrading. Also, perhaps your configure scripts could benefit from some of the new features in version 2; the changes are summarized in the file ‘NEWS’ in the Autoconf distribution.

First, make sure you have GNU m4 version 1.1 or higher installed, preferably 1.3 or higher. Versions before 1.1 have bugs that prevent them from working with Autoconf version 2. Versions 1.3 and later are much faster than earlier versions, because as of version 1.3, GNU m4 has a more efficient implementation of diversions and can freeze its internal state in a file that it can read back quickly.

13.1 Changed File Names

If you have an ‘aclocal.m4’ installed with Autoconf (as opposed to in a particular package’s source directory), you must rename it to ‘acsite.m4’. See Section 2.4 [Invoking autoconf], page 6.

If you distribute ‘install.sh’ with your package, rename it to ‘install-sh’ so make builtin rules won’t inadvertently create a file called ‘install’ from it. AC_PROG_INSTALL looks for the script under both names, but it is best to use the new name.

If you were using ‘config.h.top’ or ‘config.h.bot’, you still can, but you will have less clutter if you merge them into ‘acconfig.h’. See Section 3.4.2 [Invoking autoheader], page 16.

13.2 Changed Makefiles

   Add ‘@CFLAGS@’, ‘@CPPFLAGS@’, and ‘@LDFLAGS@’ in your ‘Makefile.in’ files, so they can

take advantage of the values of those variables in the environment when configure is run. Doing this isn’t necessary, but it’s a convenience for users.

Also add ‘@configure_input@’ in a comment to each non-‘Makefile’ input file for AC_ OUTPUT, so that the output files will contain a comment saying they were produced by configure. Automatically selecting the right comment syntax for all the kinds of files that people call AC_OUTPUT on became too much work.

Add ‘config.log’ and ‘config.cache’ to the list of files you remove in distclean targets.

If you have the following in ‘Makefile.in’:

     prefix = /usr/local
     exec_prefix = ${prefix}

you must change it to:

     prefix = @prefix@
     exec_prefix = @exec_prefix@

The old behavior of replacing those variables without ‘@’ characters around them has been removed.

13.3 Changed Macros

Many of the macros were renamed in Autoconf version 2. You can still use the old names, but the new ones are clearer, and it’s easier to find the documentation for them. See Chapter 15 [Old Macro Names], page 79, for a table showing the new names for the old macros. Use the autoupdate program to convert your ‘configure.in’ to using the new macro names. See Section 13.4 [Invoking autoupdate], page 72.

Some macros have been superseded by similar ones that do the job better, but are not call-compatible. If you get warnings about calling obsolete macros while running autoconf, you may safely ignore them, but your configure script will generally work better if you follow the advice it prints about what to replace the obsolete macros with. In particular, the mechanism for reporting the results of tests has changed. If you were using echo or AC_VERBOSE (perhaps via AC_COMPILE_CHECK), your configure script’s output will look better if you switch to AC_MSG_CHECKING and AC_MSG_RESULT. See Section 6.4 [Printing Messages], page 50. Those macros work best in conjunction with cache variables. See Section 6.3 [Caching Results], page 48.

13.4 Using autoupdate to Modernize configure

The autoupdate program updates a ‘configure.in’ file that calls Autoconf macros by their old names to use the current macro names. In version 2 of Autoconf, most of the macros were renamed to use a more uniform and descriptive naming scheme. See Section 7.2 [Macro Names], page 52, for a description of the new scheme. Although the old names still work (see Chapter 15 [Old Macro Names], page 79, for a list of the old macro names and the corresponding new names), you can make your ‘configure.in’ files more readable and make it easier to use the current Autoconf documentation if you update them to use the new macro names.

If given no arguments, autoupdate updates ‘configure.in’, backing up the original version with the suffix ‘~’ (or the value of the environment variable SIMPLE_BACKUP_SUFFIX, if that is set). If you give autoupdate an argument, it reads that file instead of ‘configure.in’ and writes the updated file to the standard output. autoupdate accepts the following options:

--help

-h Print a summary of the command line options and exit.

--macrodir=dir

-m dir Look for the Autoconf macro files in directory dir instead of the default installation directory. You can also set the AC_MACRODIR environment variable to a directory; this option overrides the environment variable.

--version

Print the version number of autoupdate and exit.

13.5 Changed Results

If you were checking the results of previous tests by examining the shell variable DEFS, you need to switch to checking the values of the cache variables for those tests. DEFS no longer exists while configure is running; it is only created when generating output files. This difference from version 1 is because properly quoting the contents of that variable turned out to be too cumbersome and inefficient to do every time AC_DEFINE is called. See Section 6.3.1 [Cache Variable Names], page 48.

For example, here is a ‘configure.in’ fragment written for Autoconf version 1:

     AC_HAVE_FUNCS(syslog)
     case "$DEFS" in
     *-DHAVE_SYSLOG*) ;;
     *) # syslog is not in the default libraries.  See if it's in some other.
       saved_LIBS="$LIBS"
       for lib in bsd socket inet; do
         AC_CHECKING(for syslog in -l$lib)
         LIBS="$saved_LIBS -l$lib"
         AC_HAVE_FUNCS(syslog)
         case "$DEFS" in
         *-DHAVE_SYSLOG*) break ;;
         *) ;;
         esac
         LIBS="$saved_LIBS"
       done ;;
     esac

Here is a way to write it for version 2:

     AC_CHECK_FUNCS(syslog)
     if test $ac_cv_func_syslog = no; then
       # syslog is not in the default libraries.  See if it's in some other.
       for lib in bsd socket inet; do
         AC_CHECK_LIB($lib, syslog, [AC_DEFINE(HAVE_SYSLOG)
           LIBS="$LIBS $lib"; break])
       done
     fi

If you were working around bugs in AC_DEFINE_UNQUOTED by adding backslashes before quotes, you need to remove them. It now works predictably, and does not treat quotes (except backquotes) specially. See Section 6.2 [Setting Output Variables], page 47.

All of the boolean shell variables set by Autoconf macros now use ‘yes’ for the true value. Most of them use ‘no’ for false, though for backward compatibility some use the empty string instead. If you were relying on a shell variable being set to something like 1 or ‘t’ for true, you need to change your tests.

13.6 Changed Macro Writing

When defining your own macros, you should now use AC_DEFUN instead of define. AC_ DEFUN automatically calls AC_PROVIDE and ensures that macros called via AC_REQUIRE do not interrupt other macros, to prevent nested ‘checking…’ messages on the screen. There’s no actual harm in continuing to use the older way, but it’s less convenient and attractive. See Section 7.1 [Macro Definitions], page 52.

You probably looked at the macros that came with Autoconf as a guide for how to do things. It would be a good idea to take a look at the new versions of them, as the style is somewhat improved and they take advantage of some new features.

If you were doing tricky things with undocumented Autoconf internals (macros, variables, diversions), check whether you need to change anything to account for changes that have been made. Perhaps you can even use an officially supported technique in version 2 instead of kludging. Or perhaps not.

To speed up your locally written feature tests, add caching to them. See whether any of your tests are of general enough usefulness to encapsulate into macros that you can share.

14 History of Autoconf

You may be wondering, Why was Autoconf originally written? How did it get into its present form? (Why does it look like gorilla spit?) If you’re not wondering, then this chapter contains no information useful to you, and you might as well skip it. If you are wondering, then let there be light…

14.1 Genesis

In June 1991 I was maintaining many of the GNU utilities for the Free Software Foundation. As they were ported to more platforms and more programs were added, the number of ‘-D’ options that users had to select in the ‘Makefile’ (around 20) became burdensome. Especially for me—I had to test each new release on a bunch of different systems. So I wrote a little shell script to guess some of the correct settings for the fileutils package, and released it as part of fileutils 2.0. That configure script worked well enough that the next month I adapted it (by hand) to create similar configure scripts for several other GNU utilities packages. Brian Berliner also adapted one of my scripts for his CVS revision control system.

Later that summer, I learned that Richard Stallman and Richard Pixley were developing similar scripts to use in the GNU compiler tools; so I adapted my configure scripts to support their evolving interface: using the file name ‘Makefile.in’ as the templates; adding ‘+srcdir’, the first option (of many); and creating ‘config.status’ files.

14.2 Exodus

As I got feedback from users, I incorporated many improvements, using Emacs to search and replace, cut and paste, similar changes in each of the scripts. As I adapted more GNU utilities packages to use configure scripts, updating them all by hand became impractical. Rich Murphey, the maintainer of the GNU graphics utilities, sent me mail saying that the configure scripts were great, and asking if I had a tool for generating them that I could send him. No, I thought, but I should! So I started to work out how to generate them. And the journey from the slavery of hand-written configure scripts to the abundance and ease of Autoconf began.

Cygnus configure, which was being developed at around that time, is table driven; it is meant to deal mainly with a discrete number of system types with a small number of mainly unguessable features (such as details of the object file format). The automatic configuration system that Brian Fox had developed for Bash takes a similar approach. For general use, it seems to me a hopeless cause to try to maintain an up-to-date database of which features each variant of each operating system has. It’s easier and more reliable to check for most features on the fly—especially on hybrid systems that people have hacked on locally or that have patches from vendors installed.

I considered using an architecture similar to that of Cygnus configure, where there is a single configure script that reads pieces of ‘configure.in’ when run. But I didn’t want to have to distribute all of the feature tests with every package, so I settled on having a different configure made from each ‘configure.in’ by a preprocessor. That approach also offered more control and flexibility.

I looked briefly into using the Metaconfig package, by Larry Wall, Harlan Stenn, and Raphael Manfredi, but I decided not to for several reasons. The Configure scripts it produces are interactive, which I find quite inconvenient; I didn’t like the ways it checked for some features (such as library functions); I didn’t know that it was still being maintained, and the Configure scripts I had seen didn’t work on many modern systems (such as System V R4 and NeXT); it wasn’t very flexible in what it could do in response to a feature’s presence or absence; I found it confusing to learn; and it was too big and complex for my needs (I didn’t realize then how much Autoconf would eventually have to grow).

I considered using Perl to generate my style of configure scripts, but decided that m4 was better suited to the job of simple textual substitutions: it gets in the way less, because output is implicit. Plus, everyone already has it. (Initially I didn’t rely on the GNU extensions to m4.) Also, some of my friends at the University of Maryland had recently been putting m4 front ends on several programs, including tvtwm, and I was interested in trying out a new language.

14.3 Leviticus

Since my configure scripts determine the system’s capabilities automatically, with no interactive user intervention, I decided to call the program that generates them Autoconfig. But with a version number tacked on, that name would be too long for old UNIX file systems, so I shortened it to Autoconf.

In the fall of 1991 I called together a group of fellow questers after the Holy Grail of portability (er, that is, alpha testers) to give me feedback as I encapsulated pieces of my handwritten scripts in m4 macros and continued to add features and improve the techniques used in the checks. Prominent among the testers were François Pinard, who came up with the idea of making an ‘autoconf’ shell script to run m4 and check for unresolved macro calls; Richard Pixley, who suggested running the compiler instead of searching the file system to find include files and symbols, for more accurate results; Karl Berry, who got Autoconf to configure TEX and added the macro index to the documentation; and Ian Taylor, who added support for creating a C header file as an alternative to putting ‘-D’ options in a ‘Makefile’, so he could use Autoconf for his UUCP package. The alpha testers cheerfully adjusted their files again and again as the names and calling conventions of the Autoconf macros changed from release to release. They all contributed many specific checks, great ideas, and bug fixes.

14.4 Numbers

In July 1992, after months of alpha testing, I released Autoconf 1.0, and converted many GNU packages to use it. I was surprised by how positive the reaction to it was. More people started using it than I could keep track of, including people working on software that wasn’t part of the GNU Project (such as TCL, FSP, and Kerberos V5). Autoconf continued to improve rapidly, as many people using the configure scripts reported problems they encountered.

Autoconf turned out to be a good torture test for m4 implementations. UNIX m4 started to dump core because of the length of the macros that Autoconf defined, and several bugs showed up in GNU m4 as well. Eventually, we realized that we needed to use some features that only GNU m4 has. 4.3BSD m4, in particular, has an impoverished set of builtin macros; the System V version is better, but still doesn’t provide everything we need.

More development occurred as people put Autoconf under more stresses (and to uses I hadn’t anticipated). Karl Berry added checks for X11. david zuhn contributed C++ support. François Pinard made it diagnose invalid arguments. Jim Blandy bravely coerced it into configuring GNU Emacs, laying the groundwork for several later improvements. Roland McGrath got it to configure the GNU C Library, wrote the autoheader script to automate the creation of C header file templates, and added a ‘--verbose’ option to configure. Noah Friedman added the ‘--macrodir’ option and AC_MACRODIR environment variable. (He also coined the term autoconfiscate to mean “adapt a software package to use Autoconf”.) Roland and Noah improved the quoting protection in AC_DEFINE and fixed many bugs, especially when I got sick of dealing with portability problems from February through June, 1993.

14.5 Deuteronomy

A long wish list for major features had accumulated, and the effect of several years of patching by various people had left some residual cruft. In April 1994, while working for Cygnus Support, I began a major revision of Autoconf. I added most of the features of the Cygnus configure that Autoconf had lacked, largely by adapting the relevant parts of Cygnus configure with the help of david zuhn and Ken Raeburn. These features include support for using ‘config.sub’, ‘config.guess’, ‘--host’, and ‘--target’; making links to files; and running configure scripts in subdirectories. Adding these features enabled Ken to convert GNU as, and Rob Savoye to convert DejaGNU, to using Autoconf.

I added more features in response to other peoples’ requests. Many people had asked for configure scripts to share the results of the checks between runs, because (particularly when configuring a large source tree, like Cygnus does) they were frustratingly slow. Mike Haertel suggested adding site-specific initialization scripts. People distributing software that had to unpack on MS-DOS asked for a way to override the ‘.in’ extension on the file names, which produced file names like ‘config.h.in’ containing two dots. Jim Avera did an extensive examination of the problems with quoting in AC_DEFINE and AC_SUBST; his insights led to significant improvements. Richard Stallman asked that compiler output be sent to ‘config.log’ instead of ‘/dev/null’, to help people debug the Emacs configure script.

I made some other changes because of my dissatisfaction with the quality of the program. I made the messages showing results of the checks less ambiguous, always printing a result. I regularized the names of the macros and cleaned up coding style inconsistencies. I added some auxiliary utilities that I had developed to help convert source code packages to use Autoconf. With the help of François Pinard, I made the macros not interrupt each others’ messages. (That feature revealed some performance bottlenecks in GNU m4, which he hastily corrected!) I reorganized the documentation around problems people want to solve. And I began a testsuite, because experience had shown that Autoconf has a pronounced tendency to regress when we change it.

Again, several alpha testers gave invaluable feedback, especially François Pinard, Jim Meyering, Karl Berry, Rob Savoye, Ken Raeburn, and Mark Eichin.

Finally, version 2.0 was ready. And there was much rejoicing. (And I have free time again. I think. Yeah, right.)

15 Old Macro Names

In version 2 of Autoconf, most of the macros were renamed to use a more uniform and descriptive naming scheme. Here are the old names of the macros that were renamed, followed by the current names of those macros. Although the old names are still accepted by the autoconf program for backward compatibility, the old names are considered obsolete. See Section 7.2 [Macro Names], page 52, for a description of the new naming scheme.

AC_ALLOCA
          AC_FUNC_ALLOCA
AC_ARG_ARRAY

removed because of limited usefulness

AC_CHAR_UNSIGNED
          AC_C_CHAR_UNSIGNED
AC_CONST  AC_C_CONST
AC_CROSS_CHECK
          AC_C_CROSS
AC_ERROR  AC_MSG_ERROR
AC_FIND_X
          AC_PATH_X
AC_FIND_XTRA
          AC_PATH_XTRA
AC_FUNC_CHECK
          AC_CHECK_FUNC
AC_GCC_TRADITIONAL
          AC_PROG_GCC_TRADITIONAL
AC_GETGROUPS_T
          AC_TYPE_GETGROUPS
AC_GETLOADAVG
          AC_FUNC_GETLOADAVG
AC_HAVE_FUNCS
          AC_CHECK_FUNCS
AC_HAVE_HEADERS
          AC_CHECK_HEADERS
AC_HAVE_POUNDBANG

AC_SYS_INTERPRETER (different calling convention)

AC_HEADER_CHECK
          AC_CHECK_HEADER
AC_HEADER_EGREP
          AC_EGREP_HEADER

AC_INLINE
          AC_C_INLINE
AC_LN_S   AC_PROG_LN_S
AC_LONG_DOUBLE
          AC_C_LONG_DOUBLE
AC_LONG_FILE_NAMES
          AC_SYS_LONG_FILE_NAMES
AC_MAJOR_HEADER
          AC_HEADER_MAJOR
AC_MINUS_C_MINUS_O
          AC_PROG_CC_C_O
AC_MMAP   AC_FUNC_MMAP
AC_MODE_T
          AC_TYPE_MODE_T
AC_OFF_T  AC_TYPE_OFF_T
AC_PID_T  AC_TYPE_PID_T
AC_PREFIX
          AC_PREFIX_PROGRAM
AC_PROGRAMS_CHECK
          AC_CHECK_PROGS
AC_PROGRAMS_PATH
          AC_PATH_PROGS
AC_PROGRAM_CHECK
          AC_CHECK_PROG
AC_PROGRAM_EGREP
          AC_EGREP_CPP
AC_PROGRAM_PATH
          AC_PATH_PROG
AC_REMOTE_TAPE

removed because of limited usefulness

AC_RESTARTABLE_SYSCALLS
          AC_SYS_RESTARTABLE_SYSCALLS
AC_RETSIGTYPE
          AC_TYPE_SIGNAL

AC_RSH removed because of limited usefulness

AC_SETVBUF_REVERSED
          AC_FUNC_SETVBUF_REVERSED

AC_SET_MAKE
          AC_PROG_MAKE_SET
AC_SIZEOF_TYPE
          AC_CHECK_SIZEOF
AC_SIZE_T
          AC_TYPE_SIZE_T
AC_STAT_MACROS_BROKEN
          AC_HEADER_STAT
AC_STDC_HEADERS
          AC_HEADER_STDC
AC_STRCOLL
          AC_FUNC_STRCOLL
AC_ST_BLKSIZE
          AC_STRUCT_ST_BLKSIZE
AC_ST_BLOCKS
          AC_STRUCT_ST_BLOCKS
AC_ST_RDEV
          AC_STRUCT_ST_RDEV
AC_SYS_SIGLIST_DECLARED
          AC_DECL_SYS_SIGLIST
AC_TEST_CPP
          AC_TRY_CPP
AC_TEST_PROGRAM
          AC_TRY_RUN
AC_TIMEZONE
          AC_STRUCT_TIMEZONE
AC_TIME_WITH_SYS_TIME
          AC_HEADER_TIME
AC_UID_T  AC_TYPE_UID_T
AC_UTIME_NULL
          AC_FUNC_UTIME_NULL
AC_VFORK  AC_FUNC_VFORK
AC_VPRINTF
          AC_FUNC_VPRINTF
AC_WAIT3  AC_FUNC_WAIT3
AC_WARN   AC_MSG_WARN
AC_WORDS_BIGENDIAN
          AC_C_BIGENDIAN

Chapter 15: Old Macro Names 82

AC_YYTEXT_POINTER
          AC_DECL_YYTEXT

Environment Variable Index 83

Environment Variable Index

This is an alphabetical list of the environment variables that Autoconf checks.

A

AC_MACRODIR . . . . . . . . . . . . . . . . . . . . . 5, 6, 7, 8, 17, 72

C

CONFIG_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 CONFIG_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 CONFIG_SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 CONFIG_SITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 CONFIG_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

S

SIMPLE_BACKUP_SUFFIX . . . . . . . . . . . . . . . . . . . . . . . . 72

Output Variable Index 84

Output Variable Index

This is an alphabetical list of the variables that Autoconf can substitute into files that it creates, typically one or more ‘Makefile’s. See Section 6.2 [Setting Output Variables], page 47, for more information on how this is done.

A

ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 AWK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

B

bindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 build . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 build_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 build_cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 build_os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 build_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 21, 37 CFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 19 configure_input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 CPP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 CXX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 CXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 20

D

datadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 DEFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

E

exec_prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 EXEEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

F

F77. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 FFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 20 FLIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

H

host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 host_alias. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 host_cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 host_os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 host_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

I

includedir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 infodir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 INSTALL_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 INSTALL_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 INSTALL_SCRIPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

K

KMEM_GROUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

L

LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 LEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 LEX_OUTPUT_ROOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 LEXLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 libexecdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 26, 27, 32 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 37 LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 localstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

M

mandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

N

NEED_SETGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

O

OBJEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 oldincludedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

P

prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 program_transform_name . . . . . . . . . . . . . . . . . . . . . . 61

R

RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

S

sbindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Output Variable Index 85

SET_MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 sharedstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 subdirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 sysconfdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

T

target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 target_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 target_cpu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 target_os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 target_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 top_srcdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

X

X_CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 X_EXTRA_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 X_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 X_PRE_LIBS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

Y

YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

Preprocessor Symbol Index 86

Preprocessor Symbol Index

This is an alphabetical list of the C preprocessor symbols that the Autoconf macros define. To work with Autoconf, C source code needs to use these names in #if directives.

__CHAR_UNSIGNED__. . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 _ALL_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 _MINIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 _POSIX_1_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 _POSIX_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 _POSIX_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

C

C_ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 CLOSEDIR_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 const . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

D

DGUX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 DIRENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

F

F77_NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . 21

G

GETGROUPS_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 GETLODAVG_PRIVILEGED . . . . . . . . . . . . . . . . . . . . . . . . 25 GETPGRP_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 gid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

H

HAVE_ALLOCA_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 HAVE_CONFIG_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 HAVE_DIRENT_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 HAVE_DOPRNT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 HAVE_function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 HAVE_GETMNTENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 HAVE_header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 HAVE_LONG_DOUBLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 HAVE_LONG_FILE_NAMES . . . . . . . . . . . . . . . . . . . . . . . . 36 HAVE_MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 HAVE_NDIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

HAVE_RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . . 36

HAVE_ST_BLKSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HAVE_ST_BLOCKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HAVE_ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HAVE_STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 HAVE_STRFTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 HAVE_STRINGIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 HAVE_SYS_DIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 HAVE_SYS_NDIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 HAVE_SYS_WAIT_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 HAVE_TM_ZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HAVE_TZNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HAVE_UNISTD_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 HAVE_UTIME_NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 HAVE_VFORK_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 HAVE_VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 HAVE_WAIT3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

I

inline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 INT_16_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

L

LONG_64_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

M

MAJOR_IN_MKDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 MAJOR_IN_SYSMACROS . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mode_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

N

NDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 NEED_MEMORY_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 NEED_SETGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 NLIST_NAME_UNION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 NLIST_STRUCT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . . . . . 20

O

off_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

P

pid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

R

RETSIGTYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

S

SELECT_TYPE_ARG1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

Preprocessor Symbol Index 87

SELECT_TYPE_ARG234 . . . . . . . . . . . . . . . . . . . . . . . . . . 26 SELECT_TYPE_ARG5. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 SETPGRP_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 SETVBUF_REVERSED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 STDC_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 SVR4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 SYS_SIGLIST_DECLARED . . . . . . . . . . . . . . . . . . . . . . . . 28 SYSDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 SYSNDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

T

TIME_WITH_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . 32 TM_IN_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

U

uid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 UMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 UMAX4_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 USG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

V

vfork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 VOID_CLOSEDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

W

WORDS_BIGENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

Y

YYTEXT_POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Macro Index 88

Macro Index

This is an alphabetical list of the Autoconf macros. To make the list easier to use, the macros are listed without their preceding ‘AC_’.

A

AIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 ARG_ARRAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 ARG_ENABLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 ARG_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 ARG_WITH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

B

BEFORE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

C

C_BIGENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C_CHAR_UNSIGNED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C_CONST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C_CROSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 C_INLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C_LONG_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C_STRINGIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 CACHE_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CACHE_LOAD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CACHE_SAVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CACHE_VAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CANONICAL_HOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 CANONICAL_SYSTEM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 CHAR_UNSIGNED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 CHECK_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CHECK_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CHECK_FUNC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 CHECK_FUNCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 CHECK_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 CHECK_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 CHECK_LIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 CHECK_PROG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CHECK_PROGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CHECK_SIZEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 CHECK_TOOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 CHECK_TYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CHECKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 COMPILE_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 CONFIG_AUX_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 CONFIG_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 CONFIG_SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 CONST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 CROSS_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 CYGWIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

D

DECL_SYS_SIGLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 DECL_YYTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 DEFINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 DEFINE_UNQUOTED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 DEFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 DIR_HEADER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 DYNIX_SEQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

E

EGREP_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 EGREP_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 EXEEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

F

F77_LIBRARY_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . 35 FIND_X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 FIND_XTRA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 FUNC_ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 FUNC_CHECK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 FUNC_CLOSEDIR_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . 25 FUNC_FNMATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 FUNC_GETLOADAVG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 FUNC_GETMNTENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_GETPGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_MEMCMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_SELECT_ARGTYPES . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_SETPGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_SETVBUF_REVERSED . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_STRFTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 FUNC_UTIME_NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FUNC_VFORK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FUNC_VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 FUNC_WAIT3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

G

GCC_TRADITIONAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 GETGROUPS_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 GETLOADAVG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

H

HAVE_FUNCS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 HAVE_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 HAVE_LIBRARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 HAVE_POUNDBANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 HEADER_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 HEADER_DIRENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 HEADER_EGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 HEADER_MAJOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 HEADER_STAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 HEADER_STDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 HEADER_SYS_WAIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 HEADER_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

I

INIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 INLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 INT_16_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 IRIX_SUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 ISC_POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

L

LANG_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 LANG_CPLUSPLUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 LANG_FORTRAN77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 LANG_RESTORE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 LANG_SAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 LINK_FILES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 LONG_64_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 LONG_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 LONG_FILE_NAMES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

M

MAJOR_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MEMORY_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 MINGW32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 MINIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MODE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MSG_CHECKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 MSG_ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 MSG_RESULT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 MSG_WARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

O

OBJEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 OBSOLETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 OFF_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 OUTPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

P

PATH_PROG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 PATH_PROGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 PATH_X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 PATH_XTRA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 PID_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PREFIX_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 PREREQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 PROG_AWK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PROG_CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PROG_CC_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PROG_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PROG_CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PROG_CXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PROG_F77_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PROG_FORTRAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PROG_GCC_TRADITIONAL . . . . . . . . . . . . . . . . . . . . . . . . 21 PROG_INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PROG_LEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PROG_LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 PROG_MAKE_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 PROG_RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 PROG_YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 PROGRAM_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PROGRAM_EGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PROGRAM_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PROGRAMS_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PROGRAMS_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 PROVIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

R

REMOTE_TAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 REPLACE_FUNCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 REQUIRE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 REQUIRE_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . . . . . . . . 80 RETSIGTYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 REVISION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 RSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

S

SCO_INTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 SEARCH_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 SET_MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 SETVBUF_REVERSED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 SIZE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 SIZEOF_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 ST_BLKSIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 ST_BLOCKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 STAT_MACROS_BROKEN . . . . . . . . . . . . . . . . . . . . . . . 32, 81 STDC_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 STRUCT_ST_BLKSIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 STRUCT_ST_BLOCKS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 STRUCT_ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 STRUCT_TIMEZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 STRUCT_TM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 SUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 SUBST_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 SYS_INTERPRETER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 SYS_LONG_FILE_NAMES . . . . . . . . . . . . . . . . . . . . . . . . . 36

SYS_RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . . . 36

SYS_SIGLIST_DECLARED . . . . . . . . . . . . . . . . . . . . . . . . 81

T

TEST_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 TEST_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 TIME_WITH_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . 81 TIMEZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 TRY_COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 TRY_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 TRY_LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 TRY_LINK_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 TRY_RUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 TYPE_GETGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_MODE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_OFF_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_PID_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_SIGNAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_SIZE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 TYPE_UID_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

U

UID_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 UNISTD_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 USG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 UTIME_NULL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

V

                                                VALIDATE_CACHED_SYSTEM_TUPLE . . . . . . . . . . . . . . . 57

VERBOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 VFORK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

W

WAIT3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 WARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 WITH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 WORDS_BIGENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

X

XENIX_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

Y

YYTEXT_POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Making configure Scripts. . . . . . . . . . . . . . . . . . . 3

2.1 Writing ‘configure.in’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Using autoscan to Create ‘configure.in’ . . . . . . . . . . . . . . . . 5 2.3 Using ifnames to List Conditionals . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Using autoconf to Create configure . . . . . . . . . . . . . . . . . . . . . 6 2.5 Using autoreconf to Update configure Scripts . . . . . . . . . . . 7

3 Initialization and Output Files. . . . . . . . . . . . . . 9

3.1 Finding configure Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.2 Creating Output Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.3 Substitutions in Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.3.1 Preset Output Variables. . . . . . . . . . . . . . . . . . . . . . . . 11 3.3.2 Build Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.3.3 Automatic Remaking . . . . . . . . . . . . . . . . . . . . . . . . . . 14

3.4 Configuration Header Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4.1 Configuration Header Templates . . . . . . . . . . . . . . . . 15 3.4.2 Using autoheader to Create ‘config.h.in’. . . . . . 16

3.5 Configuring Other Packages in Subdirectories. . . . . . . . . . . . . 17 3.6 Default Prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 Version Numbers in configure. . . . . . . . . . . . . . . . . . . . . . . . . . 18

4 Existing Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4.1 Alternative Programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.1.1 Particular Program Checks . . . . . . . . . . . . . . . . . . . . . 19 4.1.2 Generic Program and File Checks . . . . . . . . . . . . . . . 22

4.2 Library Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.3 Library Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

4.3.1 Particular Function Checks . . . . . . . . . . . . . . . . . . . . . 24 4.3.2 Generic Function Checks . . . . . . . . . . . . . . . . . . . . . . . 27

4.4 Header Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 4.4.1 Particular Header Checks . . . . . . . . . . . . . . . . . . . . . . 28 4.4.2 Generic Header Checks. . . . . . . . . . . . . . . . . . . . . . . . . 31

4.5 Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.6 Typedefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

4.6.1 Particular Typedef Checks . . . . . . . . . . . . . . . . . . . . . 33 4.6.2 Generic Typedef Checks. . . . . . . . . . . . . . . . . . . . . . . . 33

4.7 C Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.8 Fortran 77 Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . 35 4.9 System Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.10 UNIX Variants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

5 Writing Tests. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

5.1 Examining Declarations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 5.2 Examining Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.3 Examining Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 5.4 Checking Run Time Behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

5.4.1 Running Test Programs . . . . . . . . . . . . . . . . . . . . . . . . 41 5.4.2 Guidelines for Test Programs . . . . . . . . . . . . . . . . . . . 42 5.4.3 Test Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

5.5 Portable Shell Programming . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.6 Testing Values and Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.7 Multiple Cases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.8 Language Choice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

6 Results of Tests. . . . . . . . . . . . . . . . . . . . . . . . . . . 46

6.1 Defining C Preprocessor Symbols . . . . . . . . . . . . . . . . . . . . . . . . 46 6.2 Setting Output Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 6.3 Caching Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

6.3.1 Cache Variable Names . . . . . . . . . . . . . . . . . . . . . . . . . 48 6.3.2 Cache Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

6.4 Printing Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

7 Writing Macros. . . . . . . . . . . . . . . . . . . . . . . . . . . 52

7.1 Macro Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.2 Macro Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.3 Quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.4 Dependencies Between Macros . . . . . . . . . . . . . . . . . . . . . . . . . . 54

7.4.1 Prerequisite Macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.4.2 Suggested Ordering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.4.3 Obsolete Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

8 Manual Configuration . . . . . . . . . . . . . . . . . . . . . 56

8.1 Specifying the System Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 8.2 Getting the Canonical System Type . . . . . . . . . . . . . . . . . . . . . 56 8.3 System Type Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8.4 Using the System Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

9 Site Configuration . . . . . . . . . . . . . . . . . . . . . . . . 59

9.1 Working With External Software . . . . . . . . . . . . . . . . . . . . . . . . 59 9.2 Choosing Package Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 9.3 Configuring Site Details . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9.4 Transforming Program Names When Installing . . . . . . . . . . . 61

9.4.1 Transformation Options . . . . . . . . . . . . . . . . . . . . . . . . 61 9.4.2 Transformation Examples . . . . . . . . . . . . . . . . . . . . . . 62 9.4.3 Transformation Rules . . . . . . . . . . . . . . . . . . . . . . . . . . 62

9.5 Setting Site Defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

10 Running configure Scripts . . . . . . . . . . . . . . . 65

11 Recreating a Configuration . . . . . . . . . . . . . . . 66

12 Questions About Autoconf . . . . . . . . . . . . . . . 68

12.1 Distributing configure Scripts. . . . . . . . . . . . . . . . . . . . . . . . . 68 12.2 Why Require GNU m4? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 12.3 How Can I Bootstrap?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 12.4 Why Not Imake? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

13 Upgrading From Version 1 . . . . . . . . . . . . . . . 71

13.1 Changed File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 13.2 Changed Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 13.3 Changed Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 13.4 Using autoupdate to Modernize configure . . . . . . . . . . . . . 72 13.5 Changed Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 13.6 Changed Macro Writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

14 History of Autoconf . . . . . . . . . . . . . . . . . . . . . 75

14.1 Genesis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 14.2 Exodus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 14.3 Leviticus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 14.4 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 14.5 Deuteronomy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

15 Old Macro Names . . . . . . . . . . . . . . . . . . . . . . . 79

Environment Variable Index . . . . . . . . . . . . . . . . . . 83

Output Variable Index . . . . . . . . . . . . . . . . . . . . . . . 84

Preprocessor Symbol Index . . . . . . . . . . . . . . . . . . . 86

Macro Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88